### v1

Abstract
Diffusion-based generative models have emerged as powerful tools in the realm of generative modeling. Despite extensive research on denoising across various timesteps and noise levels, a conflict persists regarding the relative difficulties of the denoising tasks. While various studies argue that lower timesteps present more challenging tasks, others contend that higher timesteps are more difficult. To address this conflict, our study undertakes a comprehensive examination of task difficulties, focusing on convergence behavior and changes in relative entropy between consecutive probability distributions across timesteps. Our observational study reveals that denoising at earlier timesteps poses challenges characterized by slower convergence and higher relative entropy, indicating increased task difficulty at these lower timesteps. Building on these observations, we introduce an easy-to-hard learning scheme, drawing from curriculum learning, to enhance the training process of diffusion models. By organizing timesteps or noise levels into clusters and training models with descending orders of difficulty, we facilitate an order-aware training regime, progressing from easier to harder denoising tasks, thereby deviating from the conventional approach of training diffusion models simultaneously across all timesteps. Our approach leads to improved performance and faster convergence by leveraging the benefits of curriculum learning, while maintaining orthogonality with existing improvements in diffusion training techniques. We validate these advantages through comprehensive experiments in image generation tasks, including unconditional, class-conditional, and text-to-image generation.
Keywords:
Diffusion models Task difficulty Curriculum learning
1 Introduction
Diffusion-based generative models [18, 50, 55] have achieved significant advancements in the realm of generative tasks, demonstrating notable success across various applications such as image [8], video [17, 15], audio [27], and 3D [61, 35] generation. Specifically, their exceptional adaptability and promising performance in diverse image generation contexts, such as unconditional [22, 39], class-conditional [8], and text-conditional scenarios [2, 45], demonstrate their significant impact. Such achievements have led to a growing interest in further deepening the analysis and enhancing diffusion models.
Diffusion models [18, 55] are designed to reverse the corruption of the data through the learning process at different noise levels and over multiple time steps. Recent works have delved into the learning of diffusion models across various noise levels and timesteps, revealing different stages of diffusion models. For example, Choi et al. [5] observe that when a diffusion model performs a denoising task from large to small timestep, it first generates coarse features, then gradually generates perceptually rich content, and later refines the details. Similar observation is also identified in text-to-image diffusion models [2]. Besides this aspect, various studies have further explored the learning of diffusion models across timesteps and noise levels, elucidating their transition from denoising to generative functionalities [6], modular attributes [66], frequency characteristics [65, 31], trajectories [40], affinity [11], and variations of targets [63].
These observations have not only deepened understanding of diffusion models but have also directly contributed to improvement in diffusion models. Specifically, these insights are incorporated into their method design in various works, including loss functions [14, 5, 63], architectures [31, 2], accelerated sampling [40], representations [66], and guidance [12]. Given the tangible benefits already realized from such studies, further in-depth analysis of diffusion models across timesteps and noise levels is crucial for uncovering deeper insights and achieving unprecedented advancements in their capabilities.
In this paper, to enrich the current understanding across various timesteps and noise levels, we investigate under-explored areas within diffusion models focusing on the task difficulties of denoising. Regarding denoising task difficulties, previous works speculate that denoising tasks across timesteps have different difficulties [32, 2], yet a detailed exploration of these variances remains sparse. Moreover, there exists a notable discrepancy among studies, with works identifying larger timesteps and noise levels as more difficult [18, 14], while others argue that smaller timesteps and noise levels pose greater difficulties [22, 9, 24]. The discrepancy in difficulty across timesteps not only impedes the accurate interpretation of previous studies but also hinders the development of sophisticated training methods that properly utilize the timestep-wise variation in difficulty.
In this regard, we first analyze task difficulties in two aspects to resolve these conflicts: 1) convergence properties in the learning of denoising tasks at each timestep, and 2) the change in relative entropy between consecutive probability distributions over timesteps. In the first aspect, our analysis reveals distinct convergence behaviors across timesteps, demonstrating that models trained on larger timesteps exhibit faster convergence speeds. In the second aspect, we also observe a decrease in relative entropy as we progress to later timesteps. By integrating these observations, we confirm that denoising tasks at earlier timesteps are more difficult, as indicated by slower convergence speeds and greater changes in relative entropy.
Furthermore, building on these observations, we integrate an easy-to-hard learning scheme, a concept well-established in the curriculum learning literature [13, 26, 4, 59, 43], into the training process of diffusion models. Specifically, we organize timesteps or noise levels into clusters and train the diffusion models with descending levels of difficulty, moving from clusters categorized by higher to lower timesteps. After this curriculum process, models simultaneously learn whole timesteps as standard diffusion training [18, 55, 19] to reach the convergence point. Unlike conventional approaches where diffusion models are trained simultaneously across all timesteps, our method distinguishes itself by incorporating a sequential, order-aware training regime, reflecting an intended progression from easier to harder denoising tasks.
Building upon this foundation, our curricular approach offers several notable advantages: 1) Improved Performance and 2) Faster Convergence: By leveraging the inherent benefits of curriculum learning, our method significantly enhances the quality of generation and the speed of convergence. 3) Orthogonality with Existing Improvements: Our approach is inherently model-agnostic, ensuring broad applicability across various diffusion models. Additionally, it can be integrated with advanced diffusion training techniques, such as loss weighting [5, 14].
Finally, we empirically validate the advantages of our method by conducting comprehensive experiments across a variety of image-generation tasks. These include unconditional generation, class-conditional generation, and text-to-image generation, utilizing datasets such as FFHQ [23], ImageNet [7], and MS-COCO [33]. By integrating our curriculum learning strategy into notable architectures— DiT [42] and EDM [22]—we demonstrate the efficacy of our approach in enhancing performance, accelerating convergence speed, and maintaining compatibility with existing techniques.
2 Related Works
2.1 Diffusion Models
Diffusion models [18, 50, 55] are a group of generative models that create samples by utilizing a learned denoising process to noise. Several works have focused on improving diffusion models in various aspects, including model architectures [41, 8], sampling speed [51, 37, 34], and training objectives [14, 5, 11, 25]. These endeavors often involve investigating what diffusion models learn by dividing its process, aiming to enhance the performance of diffusion models. P2 [5] under-weights training loss functions at the clean-up stage from their observation that diffusion models learn coarse, perceptual, and removing noises at large, medium, and small timesteps. Ediff-I [2] observes that earlier sampling parts rely on conditions for generation, whereas later parts ignore the conditions. They employ multiple denoisers to address the diverse characteristics of tasks associated with different parts of the sampling process. Moreover, various works have also investigated these aspects related to timesteps [6, 66, 65, 31, 40, 11, 63] (detailed illustrations can be found in Appendix A). While our study aligns with the above works, we analyze the under-explored aspect of denoising task difficulty. Furthermore, we leverage these observations to propose a curriculum learning approach.
2.2 Denoising Difficulties on Diffusion Models
Difficulties in denoising tasks in diffusion have been referred to by various works, but this aspect is not deeply explored. Several studies just hypothesize that denoising tasks of diffusion models encompass diverse difficulties [32, 2], and there have been conflicts regarding these difficulties between previous works.
Certain studies consider denoising at larger noise levels and timesteps to be more difficult [18, 14], the focus is on the challenges associated with reconstructing data from substantial noise. For instance, Hang et al. [14] articulate that while smaller timesteps (approaching zero) may require straightforward reconstructions, such strategies become less effective at higher noise levels or in larger timesteps. Similarly, Ho et al. [18] elucidate that their approach de-emphasizes loss terms at smaller timesteps to prioritize learning on the more challenging tasks at larger timesteps, thereby enhancing sample quality. Conversely, other studies argue that earlier timesteps or lower noise levels also present significant challenges. Karras et al. [22] suggest that detecting noise at low levels is challenging due to its minimal presence. Also, Kim et al. [24] illustrate the increasing difficulty and high variance in score estimation as timesteps approach zero, disturbing stable training of models. In line with these observations, Dockhorn et al. [9] build upon insights of [24], acknowledging the complexities at near zero timesteps, where score becomes highly complex and potentially unbounded.
In this work, we aim to resolve this conflict through an in-depth analysis of convergence properties and changes in relative entropy between consecutive probability distributions across timesteps.
2.3 Curriculum Learning
Curriculum learning [3, 13, 26], inspired by human learning patterns, is a method of training models in a structured order, starting with easier tasks [43] or examples [3] and gradually increasing difficulty. As pointed out by [3], curriculum learning formulation can be viewed as a continuation method [1], which starts from a smoother objective and gradually transformed into a less smooth version until it reaches the original objective function. Through this foundation, various works have achieved improved performance and faster convergence compared to standard training based on random mini-batches sampled uniformly [13, 26, 4, 59].
Curriculum learning primarily comprises two components: a scoring function, which measures the difficulty of tasks or examples, and a pacing function, which modulates the speed of the curriculum learning process. Regarding a score function, early studies have utilized human intuition for measuring difficulty, such as the complexity of geometric shapes in images [3] or the length of sequences [56]. More recently, various works employ the model for measuring difficulty, including confidence of pre-trained models [13] and the loss of the current models [26]. For the pacing function, a predefined pacing function has been employed, which involves training using a predetermined curriculum progression [13, 62]. There are various forms of this and they can be generally represented as a function of training iteration [13, 62]. Contrary to this, there have been proposals for pacing techniques that dynamically adjust based on the loss or performance of the current model during training [29, 21].
In the diffusion model literature, curriculum learning has been utilized to organize the order of training data types based on prior knowledge of targeted generation tasks [57, 64]. Tang et al. [57] sequentially train video diffusion models with lower resolution and FPS datasets before progressing to higher resolution and FPS datasets. Similarly, Yang et al. [64] order text-to-sound generation data based on the number of events in audio clips, training diffusion models from lower to higher events datasets. In contrast, our method explores the nature of denoising task difficulty in diffusion models and proposes a curriculum learning approach that progresses from easy to hard timesteps, deviating from the standard simultaneous training of all timesteps.
Also, while consistency models [53, 52] adopt a curriculum approach to discretizing noise levels, progressively increasing the discretization steps of noise levels during training, we have distinct by exploring which noise level should be learned first and investigating the difficulty of denoising at each noise level.
3 Preliminaries
In this section, we provide the necessary background on diffusion models [18, 50, 55]. Let $\bm{x}_{0}\in\mathbb{R}^{d}$ be a sample from the data distribution $p_{0}(\bm{x})$. The forward process of diffusion models transforms data $\bm{x}_{0}$ to latent $\bm{x}_{t\in[0,T]}$ by iteratively adding Gaussian noise. This can be formulated as a stochastic differential equation (SDE) [55] as:
$\text{d}\bm{x}_{t}=f(t)\bm{x}_{t}\text{d}t+g(t)\text{d}\bm{w}_{t},$ (1)
where $f(t)$ and $g(t)$ are drift and diffusion coefficients, and $\bm{w}_{t}$ is the standard Wiener process. The Gaussian transition kernel of this SDE is formulated as:
$p_{0t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\mathbf{I}),\;\;s_{t}=\text{exp}\left(\int_{0}^{t}f(\xi)\text{d}\xi\right),\;\;\sigma_{t}=\sqrt{\int_{0}^{t}\frac{g(\xi)^{2}}{s_{\xi}^{2}}\text{d}\xi}.$ (2)
For generation, diffusion models aim to learn the corresponding reverse SDE represented as:
$\text{d}\bm{x}_{t}=\left[f(t)\bm{x}_{t}-g^{2}(t)\nabla\log p_{t}(\bm{x}_{t})\right]\text{d}\bar{t}+g(t)\text{d}\bar{\bm{w}}_{t},$ (3)
where $\bar{\bm{w}}_{t}$ and $\text{d}\bar{t}$ denote the reverse-time Wiener process and the infinitesimal reverse-time, respectively, with the actual data score $\nabla\log p_{t}(\bm{x}_{t})$. In most cases, a neural network $\epsilon_{\theta}$ having parameter $\theta$ is utilized to approximate this score function by learning the denoising tasks for each timestep $t$ from score matching loss $\mathcal{L}$ [54]:
$\mathcal{L}=\frac{1}{2}\int^{T}_{0}\mathcal{L}_{t}\text{d}t,\quad\mathcal{L}_{t}=\omega(t)\mathbb{E}_{\bm{x}_{t}\sim p_{0t}(\bm{x}_{t}|\bm{x}_{0}),\bm{x_{0}}\sim p_{0}}\left[||\epsilon_{\theta}(\bm{x}_{t},t)-\nabla\log p_{0t}(\bm{x}_{t}|\bm{x}0)||_{2}^{2}\right],$ (4)
where $\omega(t)$ is loss weights for the timestep $t$ and $p_{0t}(\bm{x}_{t}|\bm{x}_{0})$ is the transition density of $\bm{x}_{t}$ from the initial time to time $t$. This object can be interpreted as a noise-matching loss in DDPM [18], which predicts noise components in $\bm{x}_{t}$ and can be illustrated as $\int_{0}^{T}\mathbb{E}_{\bm{x}_{0}\sim p_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I})}[||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)-\epsilon||_{2}^{2}]\text{d}t$.
4 Observations
In this section, we examine the difficulties associated with learning denoising tasks across different timesteps, addressing inconsistencies in prior works regarding these difficulties. Our analysis is structured around two key aspects: 1) the convergence of loss and denoising performance across timesteps, providing insights into learning dynamics at various timestep stages in Section 4.1; and 2) the relative entropy change from $p_{t}$ to $p_{t-1}$ as a function of $t$, offering a quantitative measure of task difficulty progression over $t$ in Section 4.2. Upon integrating our findings, we establish a key conclusion: the learning difficulty for denoising tasks escalates as the timestep $t$ decreases.
4.1 Analysis on the Task Difficulty in terms of Convergence
Here, we analyze the convergence speed of loss and performance across timesteps $t$ by utilizing notable diffusion frameworks, DiT [20] and EDM [22]. The details of experimental setups in our observation can be found in Appendix B.
4.1.1 Convegence speed on loss.
First, we analyze convergence characteristics of training loss across timesteps $t$. We divided whole timesteps $[0,T]$ into 20 uniformly divided intervals and trained 20 models $\{\mathrm{M}_{i}\}_{i=1}^{20}$ where $i$-th model learns denoising tasks in $[\frac{i-1}{20}T,\frac{i}{20}T]$. During training, we tracked the loss values through iterations and plotted their convergence speed by normalizing their value in Fig. 0(a) and 0(c). As shown in the results, it is apparent that as $i$ increases towards $i=20$, the convergence accelerates in both DiT and EDM, suggesting that models learning larger timesteps can reach convergence more swiftly and reinforcing the notion that denoising tasks with larger timesteps are less difficult.
4.1.2 Convegence speed on denoising performance.
Although there is a known correlation between convergence speed of the loss function and task difficulties as indicated in prior curriculum learning literatures [26, 60], we delve deeper into the denoising performance convergence of 20 distinct models $\{\mathrm{M}_{i}\}_{i=1}^{20}$. To evaluate the performance of denoising tasks of each model, we generated samples where $\mathrm{M}_{i}$ was specifically employed for denoising within the timesteps that it was trained on, while the diffusion model learned whole timesteps handled denoising for the remaining timesteps. Then, the performance of the denoising capability of $\mathrm{M}_{i}$ can be quantitatively measured using the FID score [16], enabling us to observe the performance convergence of each model on its denoising tasks throughout the training process. Figures 0(b) and 0(d) depict this convergence. They illustrate that, similar to the loss convergence experiments, denoising performance converges more swiftly for models $\mathrm{M}_{i}$ with larger $i$ values, as observed in both DiT and EDM frameworks. These results also suggest that models trained on later timesteps, indicated by larger $i$ values, achieve faster convergence, highlighting easier task difficulty at larger timesteps.
(a) DiT (Loss)
(b) DiT (FID)
(c) EDM (Loss)
(d) EDM (FID)
Figure 1: Loss and FID scores plotted during training for each diffusion model $\mathrm{M}_{i}$ in DiT and EDM. Since the loss scale for each model is different, we show the normalized value. We observe that as $i$ increases (i.e., corresponding to larger denoising timesteps), the loss converges more rapidly, and this convergence speed correlates with that of the FID scores.
4.2 Exploration on Difficulties of Denoising Tasks
Beyond empirical convergence metrics, we also delve into analyzing the relative entropy between $p_{t}$ and $p_{t-1}$ to gain a better understanding of task difficulties from a distributional viewpoint. The training of the diffusion model implicitly involves learning the distribution of the reverse process of the corresponding SDE shown in Eq. (1). To be specific, the transition probability of the reverse process is expressed as a conditional normal distribution whose mean parameter is modeled by neural networks, and the neural networks are thereby trained to learn the dynamics of the reverse process [18]. Furthermore, the unconditional distribution of $\bm{x}_{t}$ can be obtained by marginalizing the transition densities over the prior distribution, indicating that information on the dynamics of the marginal distribution is fed to the neural networks [55].
To analyze the relationship between the dynamics of the unconditional distribution and the rate of loss convergence, we use the Kullback-Leibler (KL) divergence of $p_{t-1}$ to $p_{t}$, $D_{KL}(p_{t-1}||p_{t})$, as a quantitative measure. It is a pertinent divergence in that the training mechanism of diffusion models involves maximizing the likelihood of the reverse process. The KL divergence $D_{KL}(p_{t-1}||p_{t})$ is given by
$D_{KL}\left(p_{t-1}||p_{t}\right)=\mathbb{E}_{\bm{x}\sim p_{t-1}}\left[\log\left(\frac{p_{t-1}(\bm{x})}{p_{t}(\bm{x})}\right)\right].$ (5)
Moreover, the distribution $p_{t}$ of $\bm{x}$ at $t$ is expressed by
$p_{t}(\bm{x}_{t})=\int p_{0t}(\bm{x}_{t}|\bm{x}_{0}=\bm{y})p_{0}(\bm{y})\text{d}y=\mathbb{E}_{\bm{x}_{0}\sim p_{0}}\left[p_{0t}\left(\bm{x}_{t}|\bm{x}_{0}\right)\right].$ (6)
However, since the explicit density form of $p_{0}$ is unknown and it is computationally infeasible to estimate high-dimensional integrals in the Eqs. (5) and (6), we approximate them through unbiased estimators:
$\hat{D}_{KL}\left(p_{t-1}||p_{t}\right)=\frac{1}{M}\sum_{\begin{subarray}{c}i\in\{1,2,\cdots,M\}\\ \bm{x}_{i}\sim p_{t-1}\end{subarray}}\log\left(\frac{p_{t-1}(\bm{x}_{i})}{p_{t}(\bm{x}_{i})}\right),$ (7)
$\hat{p}_{t}(\bm{x}_{t})=\frac{1}{L}\sum_{\begin{subarray}{c}j\in\{1,2,\cdots,L\}\\ \bm{y}_{j}\sim p_{0}\end{subarray}}p_{0t}(\bm{x}_{t}|\bm{x}_{0}=\bm{y}_{j}),$ (8)
where $\hat{D}_{KL}$ and $\hat{p}_{t}$ are unbiased estimators of $D_{KL}$ and $p_{t}$, respectively, and we choose Monte-Carlo estimators for them [10, 38]. The empirical results of $D_{KL}(p_{t-1}||p_{t})$ for $64\times 64$ image data are given in Fig. 2. As seen, the relative entropy tends to decrease as $t$ increases (i.e., $D_{KL}\left(p_{s-1}||p_{s}\right)\leq D_{KL}\left(p_{t-1}||p_{t}\right)$ for $s\leq t$), which is consistent with the results in Fig. 1.
Figure 2: The KLD of $p_{t-1}$ from $p_{t}$ against denoising timestep. As the timestep increases, the dynamics decrease.
This observation may stem from the inherent low-dimensional manifold of image data. As is well-known (e.g., [48]), the image data is distributed on a relatively low-dimensional manifold with a narrow support and a highly peaked multi-modal structure. On the other hand, as Gaussian noise is iteratively added, the distribution of $\bm{x}_{t}$ approaches the independent Gaussian distribution in the ambient space. Consequently, the support of the manifold broadens and the score function becomes regular over the ambient space with increasing $t$. This nature of the unconditional distribution may cause the relative entropy from $p_{t}$ to $p_{t-1}$ to decrease with $t$, indicating that it is more difficult to accurately represent the dynamics of the reverse process at small $t$.
5 Methodology
In Section 4, we observe that denoising tasks at smaller $t$ are more difficult to learn by models. From these order of difficulties in denoising tasks, we propose the incorporation strategy of an easy-to-hard training scheme, that has demonstrated its effectiveness in curriculum literature [3, 13, 26], for improving diffusion models’ training.
5.1 Design of Curriculum Learning in Diffusion Models
As we observed in Section 4, difficulties in denoising tasks increase as $t$ gets smaller. To utilize this in our training strategy for diffusion models, we adopt an easy-to-hard curriculum learning approach by partitioning the entire range of timesteps into $N$ clusters, denoted as $\{C_{i}\}_{i=1}^{N}$, where each cluster $C_{i}$ spans an interval $[l_{i},r_{i}]$, ensuring $l_{i}<r_{i}$ and continuity with $r_{i}=l_{i+1}$ for adjacent clusters. We then arrange for the denoising tasks to become progressively harder as we move from cluster $N$ to cluster $1$, ensuring $l_{i}<l_{i+1}$, with $l_{1}=0$ and $r_{N}=T$. During training, we introduce these clusters sequentially, starting from the least challenging (the $N$-th cluster) and advancing towards the most challenging (the $1$-st cluster), through $N$ distinct phases. Specifically, in the $n$-th stage, we jointly train the model on clusters $\{C_{N},C_{N-1},\ldots,C_{N-(n-1)}\}$. After this $N$-phase of curriculum learning, the model continues to learn across the entire range of timesteps, progressively converging towards a point of stability.
The next consideration involves determining the boundaries for each cluster, $l_{i}$ and $r_{i}$. A straightforward approach is to uniformly divide the entire timestep interval $[0,T]$ as $C_{i}=[\frac{(i-1)\cdot T}{N},\frac{i\cdot T}{N}]$ for $i=1,2,\ldots,N$. However, this method does not account for variations in noise levels across different timesteps. Therefore, to address this issue more effectively, we adopt an SNR-based interval clustering technique as used in [11], which aligns the clustering with the actual changes in noise levels, potentially enhancing curriculum learning adaptability to varying noise conditions.
In contrast, for EDM [23] which operates based on the noise level $\sigma$ rather than the timestep $t$, and where $\sigma$ is sampled from a log-normal distribution such that $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ during training, our clustering strategy for timesteps cannot be directly transposed. Given the log-normal distribution of $\sigma$, dividing it directly is impractical because $\sigma$ can extend over a wide range of values. Instead, we adapt our clustering approach to suit the log-normal characteristics by defining noise level clusters $C_{i}$. Specifically, we delineate $C_{i}=[\mathrm{\Phi}^{-1}(\frac{i-1}{N}),\mathrm{\Phi}^{-1}(\frac{i}{N})]$, where $\mathrm{\Phi}^{-1}$ is the inverse cumulative distribution function (quantile function) of the Gaussian distribution $\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$. This method segments the noise levels into intervals by reflecting their probabilistic distribution.
Algorithm 1 Curriculum Learning Input: Curriculum $\{C_{i}\}_{i=1}^{N}$, Pacing function $g$, Maximum patience $\tau_{\text{max}}$, Loss function $f$, Curriculum index $I_{\text{cur}}=N$, Best loss $L_{\text{best}}=\infty$, Model $M_{\theta}$ while $I_{\text{cur}}>0$ do     # Mini-batch sampling     $X\sim C_{I_{\text{cur}}}$     # Calculate Loss     $L_{\text{cur}}=f(M_{\theta}(X))$     # Update model     $\theta$ = $\theta$ $\nabla_{\theta}{L_{\text{cur}}}$     # Pacing function     $\tau_{\text{cur}},I_{\text{next}}=g(L_{\text{cur}},L_{\text{best}},\tau_{\text{cur}},\tau_{\text{max}},I_{\text{cur}})$     # Update curriculum     if $I_{\text{cur}}\neq I_{\text{next}}$ then        $I_{\text{cur}}=I_{\text{next}}$        $L_{\text{best}}=\infty$    # Update best loss     else if $L_{\text{cur}}<L_{\text{best}}$ then        $L_{\text{best}}=L_{\text{cur}}$     end if end while
Algorithm 2 Pacing Function Input: Current loss $L_{\text{cur}}$, Best loss $L_{\text{best}}$, Current patience $\tau_{\text{cur}}$, Maximum patience $\tau_{\text{max}}$, Current curriculum index $I_{\text{cur}}$ Output: Updated patience, Updated curriculum index # Reset patience if $L_{\text{cur}}<L_{\text{best}}$ then     return 0, $I_{\text{cur}}$ else     # Proceed to next curriculum     if $\tau_{\text{cur}}+1>\tau_{\text{max}}$ then        return 0, $I_{\text{cur}}-1$    # Increase patience     else        return $p_{\text{cur}}+1,I_{\text{cur}}$     end if end if
5.2 Pacing Strategy of Curriculum
To effectively train the diffusion model according to the provided curriculum design, it is crucial to define a suitable pacing function for determining the transition of each $N$ distinct curriculum stage. Training for a fixed number of iterations for each curriculum is the simplest implementation (We also contain this method in experiments as ‘NaiveCL’ in Section 6.2). However, the convergence rate of each curriculum varies significantly, as demonstrated in Fig. 1. Hence, we propose adopting an adaptive number of iterations for each curriculum, akin to the varied exponential pacing approach explored by Hacohen et al. [13]. The proposed pacing function considers the loss of a mini-batch and maximum patience. If the loss does not improve within the maximum patience iterations, the current curriculum is terminated, and the subsequent curriculum is initiated. Here, the maximum patience is a fixed hyper-parameter and the detailed process is outlined in Algorithm 2. The overall curriculum learning procedure is presented in Algorithm 1.
6 Experimental Results
In this section, we present experimental results to validate the effectiveness of our method. The advantages of our curriculum method, 1) Improved Performance, 2) Faster Convergence, and 3) Orthogonality with Existing Improvements, are validated in this section. To begin, we outline our experimental setups in Section 6.1. Then, we provide the results of the comparative evaluation in Section 6.2, showing that our curriculum approach significantly improves the quality of generated samples compared to the baseline. Finally, analyses of our method are illustrated in Section 6.3 to deeply understand the effectiveness of our method.
6.1 Experimental Setup
Here, we provide experimental setups concisely due to space constraints. Detailed setups are presented in Appendix C.
6.1.1 Evaluation protocols.
For our comprehensive evaluation of various methods, we employed three distinct image-generation tasks: 1) Unconditional generation: We utilized the FFHQ dataset [23], which comprises 70,000 training images of human faces. 2) Class-conditional generation: For this task, we employed two datasets: CIFAR-10 [28], which includes 60,000 training images, and ImageNet [7], containing 1,281,167 training images. In this setup, we applied classifier-free guidance [19]. 3) Text-to-Image generation: We used the MS-COCO dataset [33] , which contains 82,783 training images and 40,504 validation images. Each image in this dataset is accompanied by five text captions, facilitating the text-to-image generation task.
6.1.2 Evaluation metrics.
To evaluate the performance of models, we utilized three metrics: FID [16], IS [49], and Precision/Recall [30]. Specifically, we applied FID and IS to measure sample quality, while Precision is used to assess quality further and Recall was utilized to evaluate the diversity of the generated samples in ImageNet setup. In other datasets, we employed FID to evaluate sample quality.
6.1.3 Target models.
We employed two exemplary diffusion architectures for experiments: DiT [42], which integrates latent diffusion models [46] with Transformer architectures [58], and EDM [22], which focuses on pixel-level diffusion utilizing UNet-based architectures [47]. For the text-to-image generation, we incorporated a CLIP text encoder [44] as described in DTR [41].
6.2 Comparative Results
In this section, we assess the effectiveness of our curriculum-based training approach. For a thorough comparison, we examine three distinct training variants, with further details provided in Appendix C: To achieve this, we compare three variants of training: 1) Vanilla: This term refers to diffusion models trained using conventional methods without any curriculum learning strategies; 2) NaiveCL: In this variant, we incorporate a basic curriculum learning strategy, which simply repeats the same number of iterations for each stage across an $N$-stage process and does not employ SNR-based clustering; 3) Ours: This denotes our proposed curriculum approach, which is designed to enhance the training process of diffusion models by systematically structuring the learning stages.
6.2.1 Quantitative evaluation.
As shown in Table 1, our approach was validated on various architectures—DiT [42] and EDM [22]— and tasks including unconditional, class-conditional, and text-to-image generation. To further contextualize our method, we evaluated the performance of NaiveCL, which does not leverage the SNR-based clustering and pacing function proposed in our approach. Our observations indicate that NaiveCL fails to consistently achieve improved performance, as it lacks the capability to incorporate task difficulty. In contrast, our method explicitly integrates the difficulty of denoising tasks, as detailed in Section 4. Consequently, our approach consistently demonstrates significant performance enhancements across all metrics on four datasets: FFHQ [23], ImageNet [7], CIFAR-10 [28], and MS-COCO [33], demonstrating that our approach works well data- and model-agnostic.
Latent-Level Diffusion Model.
Model FFHQ 256$\times$256 ImageNet 256$\times$256 COCO 256$\times$256
[…]
DiT (Vanilla) 10.49 11.18 146.95 0.75 0.47 7.62
DiT + NaiveCL 7.95 11.90 151.66 0.75 0.47 7.71
DiT + Ours 7.55 8.18 186.37 0.79 0.47 7.51
Pixel-Level Diffusion Model.
Model FFHQ 64$\times$64 CIFAR10 32$\times$32
[…]
EDM (Vanilla) 2.93 2.67
EDM + NaiveCL 3.13 2.88
EDM + Ours 2.71 2.44
Table 1: We evaluated unconditional image generation on FFHQ with DiT-B and EDM, class-conditional image generation on ImageNet and CIFAR10 with DiT-L and EDM, respectively, and text-conditional image generation on MS-COCO with DiT-B. Note that our curriculum learning for diffusion models improves substantial performance without any additional parameters.
6.2.2 Qualitative evaluation.
Due to space constraints, we illustrate a detailed collection of generated examples in Appendix D. In summary, our curriculum methodology demonstrates a notable enhancement in the quality of the images produced, especially when compared to naive training.
6.3 Analysis
To elucidate the effectiveness of our curriculum approach, we hereby present a series of analytical studies. All the analysis is conducted by using the DiT-B model on the ImageNet dataset.
6.3.1 Effects of $N$ and $\tau$.
Figure 3: Ablation study on $N$ and $\tau$. We use DiT-B on ImageNet 256$\times$256.
We examined the robustness of the proposed curriculum training with respect to hyper-parameters: the number of clusters $N$ and the maximum patience $\tau$. As shown in Fig. 3  our method consistently outperforms the vanilla model, and the best result is observed at $N=20,\tau=200$. It shows that as $\tau$ increases, it may lead to overtraining due to excessive iterations for each task, whereas as $\tau$ decreases, curriculums may not be sufficiently trained. Furthermore, when the entire range of timesteps is finely partitioned (i.e., with an increase in $N$), each cluster becomes excessively granular, resulting in suboptimal performance. Conversely, with a decrease in $N$, tasks that should be in distinct clusters are learned together, forming a coarser cluster, which also leads to suboptimal outcomes. We utilized the optimal hyper-parameters for all subsequent experiments.
6.3.2 Effects of Curriculum Design.
Class-Conditional ImageNet 256$\times$256.
[…]
(a) Vanilla 30.27 60.06 0.55 0.52
(b1) + anti-curriculum + uniform 31.12 62.80 0.55 0.53
(b2) + anti-curriculum + SNR 27.74 68.10 0.58 0.52
(c1) + curriculum + uniform 25.01 71.99 0.58 0.53
(c2) + curriculum + SNR 22.96 75.98 0.62 0.52
Table 2: Comparative results on various curriculum designs.
In our curriculum design, we initially partitioned the entire set of timesteps into $N$ clusters using SNR-based clustering, organizing the curriculum from easy to hard clusters. To thoroughly assess the impact of each component, we conducted the ablation study as shown in Table. 2. Firstly, we investigated the effect of curriculum learning via comparison with an anti-curriculum approach [13], which progresses from hard to easy clusters. While both training methods that consider task order (b2, c2) appear to enhance performance compared to vanilla training (a), anti-curriculum training cannot consistently guarantee performance improvement. To specifically assess the impact of anti-curriculum learning, we evaluate the performance of anti-curriculum learning without SNR-based clustering (b1), resulting in a performance decrease compared to vanilla. This can be attributed to the deviation from the observed appropriate task sequence, with the improvement seen in (b2) attributed to SNR-based clustering. In contrast, the proposed curriculum learning method (c1, c2) consistently exhibited performance improvement even with the uniformly partitioned clusters. Besides, with the finding that utilizing SNR-clustering was more effective, the clustering with the actual changes in noise levels enhanced the curriculum learning adaptability. We believe that by embracing more advanced clustering methods, we could gain further performance improvements.
6.3.3 Visualization of curriculum.
Figure 4: We visualized the curriculum transition and the corresponding loss across iterations ($N=20$). To make the loss graph more easily readable, the y-axis was truncated to 1.0.
Figure 4: We visualized the curriculum transition and the corresponding loss across iterations ($N=20$). To make the loss graph more easily readable, the y-axis was truncated to 1.0.
Figure 5: The models trained using the proposed curriculum learning approach demonstrate faster convergence compared to vanilla models, irrespective of evaluation metrics.
To gain deeper insights into the functioning of our curriculum pacing, we plotted the loss metrics against the curriculum phases, as illustrated in Fig. 5. During the curriculum training, tasks progressively transition from the easiest to the most challenging, with varying amounts of iterations for each task based on the pacing function. The training loss decreased during each curriculum phase but increased after curriculum changes via the pacing function due to the inclusion of a newly added task in the updated curriculum. In addition, as $\tau$ increases, it is confirmed that the curriculum phase slowly changes.
6.3.4 Analysis on convergence speed.
As demonstrated in previous works [3, 13], the adoption of curriculum learning can lead to faster convergence in model performance. To illustrate the efficacy of our approach in this regard, we plotted the FID, IS, precision, and recall calculated over 10,000 samples across the training iterations, as depicted in Fig. 5. We observed the models trained through the proposed curriculum learning method converge faster than vanilla models, regardless of evaluation metrics. Note that our approach does not need any additional parameters and training iterations.
6.3.5 Robustness on various sizes of models
To verify the generalizability of our method across different model sizes, we evaluated the performance gains achieved using our curriculum learning approach on various scales of the DiT model: DiT-S (small), DiT-B (base), and DiT-L (large). Table. 4 shows that the proposed curriculum learning for diffusion model improves the performance regardless of the model size. Moreover, it is notable that larger models exhibit a more substantial performance enhancement: DiT-S has improved by 8% in terms of FID, while DiT-B and DiT-L have shown improvements of 24% and 27%, respectively. These findings validate the efficacy of our curriculum approach across a diverse range of model sizes, underscoring its robust generalizability.
6.3.6 Additional experimental results
Due to limited space, we present additional experimental results in Appendix E. These results also support the effectiveness of our method, emphasizing the importance of curriculum approaches in diffusion training.
6.4 Orthogonality of Our Curriculum Approach
Lastly, we illustrate the seamless integration of our method with sophisticated training techniques such as DTR [41] and MinSNR [14]. Initially, we observed that each sophisticated method yields a superior performance compared to the vanilla method. Meanwhile, as shown in Table. 4, the performance is significantly enhanced when we apply the proposed curriculum learning. Consequently, the curriculum approach proves to be compatible with previous promising methods such as loss weighting (MinSNR) and architectural enhancements (DTR), demonstrating our orthogonality with recent diffusion techniques.
Table 3: Note that the curriculum learning achieves consistent improvements across the model sizes.
[…]
Table 3: Note that the curriculum learning achieves consistent improvements across the model sizes.
Table 4: Note that the curriculum learning is compatible with the previous works such as loss weighting (MinSNR) and architecture (DTR) study which, specified the multi-task learning for diffusion model.
7 Conclusion
In this study, we tackle the challenge of denoising task difficulty within the diffusion model framework and introduce a novel task difficulty-based curriculum learning approach. To the best of our knowledge, we are the first to define task difficulty by considering both the convergence rates of loss and performance metrics. Moreover, in terms of data distribution analysis, we observe a reduction in relative entropy between consecutive probability distributions as timesteps progress. We believe that these observations might help reorganize the conflicts of previous works regarding denoising task difficulties. Building upon these insights, we propose a curriculum learning framework for the diffusion model, comprising curriculum design and pacing strategies. Our experimental results convincingly demonstrate the efficacy of our approach across diverse diffusion model designs, datasets, and tasks. From these results, we emphasize that considering an order of learning denoising tasks is also a potential direction to improve the training of diffusion models. In future research, for further enhancements, more advanced curriculum learning strategies such as self-pacing can be elaborated.
A. Extended Related Work
In this section, we review works related to analyzing diffusion model behaviors in each timestep but not covered in detail in Section 2.1. Deja et al. [6] explore denoising during the backward diffusion process and observe that transition from denoising to generation exists in the backward process. Go et al. [11] investigate the affinity between denoising tasks, showing that temporal proximal denoising tasks exhibit higher task affinity. Then, they also observe that simultaneously learning all denoising tasks by one model suffers from negative transfer. They can achieve better performance than standard diffusion training by alleviating negative transfer. Lee et al. [31] analyze frequency characteristics according to timesteps and observe that high-frequency components are lost as timesteps increase. From this observation, they propose a multi-architecture multi-experts diffusion model, which utilizes multiple denoiser models specialized in each timestep interval but utilizes a transformer-like model as the timestep increases. From observations that smaller and larger models produce similar latent noise, Pan et al. [40] propose T-Stitch, which leverages a pre-trained smaller model at the beginning of the backward process to accelerate the sampling speed. Xu et al. [63] investigate the average trace-of-covariance of training targets according to timesteps, showing that it peaks in the intermediate timesteps, causing unstable training targets. For more stable training targets, they utilize weighted conditional scores with a reference batch.
B. Detailed Experimental Setups for Observation
In Section 4, we examined the difficulty of denoising tasks in terms of convergence with various models $\{\mathrm{M}\}_{i=1}^{20}$, which are trained within specific timesteps $[\frac{i-1}{20}T,\frac{i}{20}T]$. For the DiT architecture, we employed the DiT-B/2, whereas for EDM, we used the DDPM++ architecture. Both DiT and EDM models were trained on the FFHQ dataset, with a batch size of 256, for approximately 20,000 and 4,000 iterations (equivalent to processing 1 million images), respectively. This training was conducted until both loss and performance converged.
To evaluate the performance of diffusion models through the FID score of the generated images, performing a recursive denoising task from $T$ to zero is necessary, complicating the assessment using only $\mathrm{M}_{i}$. Following [11], we generated samples where $\mathrm{M}_{i}$ was specifically utilized for denoising within its trained intervals. At the same time, the diffusion model is responsible for the denoising tasks across the entire range of timesteps. For this evaluation, we sampled 10,000 images using a DDIM sampler over 250 steps for DiT, and an Euler solver over 40 steps for the other models.
Vanilla
Naive
Ours
Fig. A: Qualitative comparison between vanilla, naive curriculum, and ours on FFHQ dataset.
Vanilla
Naive
Ours
  Ostrich Clumber Arctic wolf Puffer American eagle Hummingbird Ridgeback
Fig. B: Qualitative comparison between vanilla, naive curriculum, and ours on ImageNet dataset.
C. Experiment Details
For training diffusion models, we utilized the AdamW optimizer [36] with a constant learning rate of 0.0001, and weight decay was not applied. The exponential moving average (EMA) to the model’s weights was used to stabilize the training and the decay ratio was set to 0.9999. The batch size was set to 256, and we augmented the training data by a horizontal flip. While the diffusion timestep $T$ was configured as 1,000 for all experiments, we trained for 100K iterations for the FFHQ dataset [23], and 400K iterations for the ImageNet dataset [7] and MS-COCO dataset [33]. The number of clusters $N$ was 20 unless otherwise specified. The maximum patience $\tau$ was varied across model sizes: it was set at 200 for DiT-S/2, DiT-B/2, and EDM, and 400 for DiT-L/2. EDM was trained using fp16, while the other models were trained using fp32.
Unless otherwise mentioned, we sampled 50K samples for evaluation. In tasks involving conditional generation, including class-conditional image generation (e.g. CIFAR-10, ImageNet) and text-to-image conversion (e.g. MS-COCO), we adapted the classifier-free guidance [19] with a guidance scale of 1.5.
Vanilla
Naive
Ours
“A man is on a path riding a horse.”
“Tower of a brick building with a clock on the front.”
“A dog leaps through the air as it catches a frisbee.”
“A person in white and red snow suit skiing on a slope.”
“A group of young people playing a game of soccer.”
“A young boy is smiling and he has food around him on a table.”
“a small boy with a hat is standing on a surf board.”
Fig. C: Qualitative comparison between vanilla, naive curriculum, and ours on MS-COCO dataset.
D. Qualitative Results
7.0.1 Qualitative Evaluation on the FFHQ Dataset.
Figure A presents a qualitative analysis of the performance in unconditional facial image synthesis among the vanilla, the naive curriculum approach, and our method. Our approach demonstrates superior ability in generating realistic images.
7.0.2 Qualitative Analysis on the ImageNet Dataset.
In the conditional image synthesis, we exhibit the outcomes generated by the vanilla, the naive curriculum strategy, and our proposed method. Figure B clearly shows that our methodology surpasses the competing approaches in terms of quality.
7.0.3 Qualitative Assessment on the MS-COCO Dataset.
To further substantiate the effectiveness of our proposed technique, we conduct a qualitative comparison of the results in the text-to-image generation task among the vanilla, the naive curriculum method, and our own approach, as depicted in Fig. C.
E. Further Experimental Results
E.1. Convergence Speed across Model Size
Fig. D: We observed an increase in convergence speed across various model sizes when the proposed curriculum learning approach was applied.
By leveraging the advantages of curriculum learning in diffusion training, our method offers faster convergence than vanilla training. To further investigate this aspect, we measured FID-10K through training iterations for DiT-S and DiT-L. Figure D describes the results, showing that our curriculum approach achieves faster convergence in both models. These results also support the effectiveness of our method.
E.2. Robustness on Noise Schedule
Class-Conditional ImageNet 256$\times$256.
[…]
cosine Vanilla 30.27 60.06 0.55 0.52
Ours 22.22 75.98 0.62 0.52
linear Vanilla 16.99 83.62 0.68 0.53
Ours 16.03 87.66 0.69 0.53
Table A: Ablation study on noise scheduler. Note that our approach improves performance consistently across each scheduler.
For more comprehensive ablation study, we also trained the diffusion model with different noise schedule. In contrast to cosine scheduling, the $\beta_{t}$ is set by uniformly dividing the interval $[0.0001,0.02]$, and the $C_{i}$ are obtained corresponding to SNR on a linear schedule. As shown in Table. A, our approach improves the performance with cosine and linear noise schedulers both.
F. Negative Social Impacts
Generative models, such as diffusion models, have the potential to significantly impact society, particularly through DeepFake applications and the use of biased datasets. One primary concern is the possibility for these models to amplify misinformation, which can erode trust in visual media. Additionally, if these models are trained on biased or deliberately altered content, they may unintentionally perpetuate and intensify existing social biases. This situation may result in the dissemination of incorrect information and the manipulation of public opinion.

### v2

@@
-
+Denoising Task Difficulty-based Curriculum for Training Diffusion Models
 Abstract
@@
 Building on these observations, we introduce an easy-to-hard learning scheme, drawing from curriculum learning, to enhance the training process of diffusion models.
-By organizing timesteps or noise levels into clusters and training models with descending orders of difficulty, we facilitate an order-aware training regime, progressing from easier to harder denoising tasks, thereby deviating from the conventional approach of training diffusion models simultaneously across all timesteps.
-Our approach leads to improved performance and faster convergence by leveraging the benefits of curriculum learning, while maintaining orthogonality with existing improvements in diffusion training techniques.
+By organizing timesteps or noise levels into clusters and training models with ascending orders of difficulty, we facilitate an order-aware training regime, progressing from easier to harder denoising tasks, thereby deviating from the conventional approach of training diffusion models simultaneously across all timesteps.
+Our approach leads to improved performance and faster convergence by leveraging benefits of curriculum learning, while maintaining orthogonality with existing improvements in diffusion training techniques.
 We validate these advantages through comprehensive experiments in image generation tasks, including unconditional, class-conditional, and text-to-image generation.
-Keywords:
-Diffusion models Task difficulty Curriculum learning
 1 Introduction
-Diffusion-based generative models [18, 50, 55] have achieved significant advancements in the realm of generative tasks, demonstrating notable success across various applications such as image [8], video [17, 15], audio [27], and 3D [61, 35] generation.
+Diffusion-based generative models [18, 52, 57] have achieved significant advancements in the realm of generative tasks, demonstrating notable success across various fields such as image [8], video [17, 15], and 3D [62, 35] generation.
 Specifically, their exceptional adaptability and promising performance in diverse image generation contexts, such as unconditional [22, 40], class-conditional [8], and text-conditional scenarios [2, 47], demonstrate their significant impact.
@@
 Specifically, these insights are incorporated into their method design in various works, including loss functions [14, 5, 64], architectures [31, 2], accelerated sampling [41], representations [67], and guidance [12].
-Given the tangible benefits already realized from such studies, further in-depth analysis of diffusion models across timesteps and noise levels is crucial for uncovering deeper insights and achieving unprecedented advancements in their capabilities.
+Given the tangible benefits already realized from such studies, further in-depth analysis of diffusion models across timesteps and noise levels is crucial for uncovering insights and achieving unprecedented advancements in their capabilities.
 In this paper, to enrich the current understanding across various timesteps and noise levels, we investigate under-explored areas within diffusion models focusing on the task difficulties of denoising.
 Regarding denoising task difficulties, previous works speculate that denoising tasks across timesteps have different difficulties [32, 2], yet a detailed exploration of these variances remains sparse.
-Moreover, there exists a notable discrepancy among studies, with works identifying larger timesteps and noise levels as more difficult [18, 14], while others argue that smaller timesteps and noise levels pose greater difficulties [22, 9, 24].
+Moreover, there exists a notable discrepancy among studies, with works identifying larger timesteps as more difficult [18, 14], while others argue that smaller timesteps pose greater difficulties [22, 9, 25].
 The discrepancy in difficulty across timesteps not only impedes the accurate interpretation of previous studies but also hinders the development of sophisticated training methods that properly utilize the timestep-wise variation in difficulty.
 In this regard, we first analyze task difficulties in two aspects to resolve these conflicts: 1) convergence properties in the learning of denoising tasks at each timestep, and 2) the change in relative entropy between consecutive probability distributions over timesteps.
-In the first aspect, our analysis reveals distinct convergence behaviors across timesteps, demonstrating that models trained on larger timesteps exhibit faster convergence speeds.
+In the first aspect, our analysis reveals distinct convergence behaviors across timesteps, demonstrating that models trained on larger timesteps exhibit faster convergence.
 In the second aspect, we also observe a decrease in relative entropy as we progress to later timesteps.
-By integrating these observations, we confirm that denoising tasks at earlier timesteps are more difficult, as indicated by slower convergence speeds and greater changes in relative entropy.
+By integrating these, we confirm that denoising tasks at earlier timesteps are more difficult, indicated by slower convergence speeds and greater changes in relative entropy.
 Furthermore, building on these observations, we integrate an easy-to-hard learning scheme, a concept well-established in the curriculum learning literature [13, 27, 4, 61, 45], into the training process of diffusion models.
-Specifically, we organize timesteps or noise levels into clusters and train the diffusion models with descending levels of difficulty, moving from clusters categorized by higher to lower timesteps.
+Specifically, we organize timesteps or noise levels into clusters and train the diffusion models with ascending levels of difficulty, moving from clusters categorized by higher to lower timesteps.
 After this curriculum process, models simultaneously learn whole timesteps as standard diffusion training [18, 57, 19] to reach the convergence point.
@@
 3) Orthogonality with Existing Improvements: Our approach is inherently model-agnostic, ensuring broad applicability across various diffusion models.
-Additionally, it can be integrated with advanced diffusion training techniques, such as loss weighting [5, 14].
+Additionally, it can be integrated with advanced diffusion training techniques, such as loss weighting [5, 14, 11, 23].
 Finally, we empirically validate the advantages of our method by conducting comprehensive experiments across a variety of image-generation tasks.
 These include unconditional generation, class-conditional generation, and text-to-image generation, utilizing datasets such as FFHQ [24], ImageNet [7], and MS-COCO [33].
-By integrating our curriculum learning strategy into notable architectures— DiT [42] and EDM [22]—we demonstrate the efficacy of our approach in enhancing performance, accelerating convergence speed, and maintaining compatibility with existing techniques.
+By integrating our curriculum learning strategy into architectures— DiT [44], EDM [22], and SiT [38]—we demonstrate the efficacy of our approach in enhancing performance, accelerating convergence speed, and maintaining compatibility with existing techniques.
 2 Related Works
@@
 Diffusion models [18, 52, 57] are a group of generative models that create samples by utilizing a learned denoising process to noise.
-Several works have focused on improving diffusion models in various aspects, including model architectures [41, 8], sampling speed [51, 37, 34], and training objectives [14, 5, 11, 25].
+Several works have focused on improving diffusion models in various aspects, including model architectures [43, 8, 42], sampling speed [53, 37, 34], training objectives [14, 5, 11, 26, 38].
 These endeavors often involve investigating what diffusion models learn by dividing its process, aiming to enhance the performance of diffusion models.
@@
 Difficulties in denoising tasks in diffusion have been referred to by various works, but this aspect is not deeply explored.
-Several studies just hypothesize that denoising tasks of diffusion models encompass diverse difficulties [32, 2], and there have been conflicts regarding these difficulties between previous works.
+Several studies hypothesize that denoising tasks in diffusion encompass diverse difficulties [32, 2], and there have been conflicts regarding these difficulties between previous works.
 Certain studies consider denoising at larger noise levels and timesteps to be more difficult [18, 14], the focus is on the challenges associated with reconstructing data from substantial noise.
@@
 Through this foundation, various works have achieved improved performance and faster convergence compared to standard training based on random mini-batches sampled uniformly [13, 27, 4, 61].
-Curriculum learning primarily comprises two components: a scoring function, which measures the difficulty of tasks or examples, and a pacing function, which modulates the speed of the curriculum learning process.
-Regarding a score function, early studies have utilized human intuition for measuring difficulty, such as the complexity of geometric shapes in images [3] or the length of sequences [56].
-More recently, various works employ the model for measuring difficulty, including confidence of pre-trained models [13] and the loss of the current models [26].
+Curriculum learning primarily comprises two components: a curriculum scoring function, measuring the difficulty of tasks or examples, and a pacing function, modulating the speed of the curriculum progress.
+Regarding a curriculum score function, early studies have utilized human intuition for measuring difficulty, such as the complexity of geometric shapes in images [3] or the length of sequences [58].
+Recently, various works employ models to measure difficulty, including confidence of pre-trained models [13] and the loss of the current models [27].
 For the pacing function, a predefined pacing function has been employed, which involves training using a predetermined curriculum progression [13, 63].
 There are various forms of this and they can be generally represented as a function of training iteration [13, 63].
-Contrary to this, there have been proposals for pacing techniques that dynamically adjust based on the loss or performance of the current model during training [29, 21].
+Contrary to this, there have been proposals for pacing techniques dynamically adjusting based on the loss or performance of the current model during training [29, 21].
 In the diffusion model literature, curriculum learning has been utilized to organize the order of training data types based on prior knowledge of targeted generation tasks [59, 65].
@@
 The forward process of diffusion models transforms data $\bm{x}_{0}$ to latent $\bm{x}_{t\in[0,T]}$ by iteratively adding Gaussian noise.
-This can be formulated as a stochastic differential equation (SDE) [55] as:
-$\text{d}\bm{x}_{t}=f(t)\bm{x}_{t}\text{d}t+g(t)\text{d}\bm{w}_{t},$ (1)
-where $f(t)$ and $g(t)$ are drift and diffusion coefficients, and $\bm{w}_{t}$ is the standard Wiener process. The Gaussian transition kernel of this SDE is formulated as:
+This can be formulated as a stochastic differential equation (SDE) [57] as $\text{d}\bm{x}_{t}=f(t)\bm{x}_{t}\text{d}t+g(t)\text{d}\bm{w}_{t}$, where $f(t)$ and $g(t)$ are drift and diffusion coefficients, and $\bm{w}_{t}$ is the standard Wiener process.
+The Gaussian transition kernel of this SDE is formulated as:
 $p_{0t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\mathbf{I}),\;\;s_{t}=\text{exp}\left(\int_{0}^{t}f(\xi)\text{d}\xi\right),\;\;\sigma_{t}=\sqrt{\int_{0}^{t}\frac{g(\xi)^{2}}{s_{\xi}^{2}}\text{d}\xi}.$ (1)
@@
 $\mathcal{L}=\frac{1}{2}\int^{T}_{0}\mathcal{L}_{t}\text{d}t,\quad\mathcal{L}_{t}=\omega(t)\mathbb{E}_{\bm{x}_{t}\sim p_{0t}(\bm{x}_{t}|\bm{x}_{0}),\bm{x_{0}}\sim p_{0}}\left[||\epsilon_{\theta}(\bm{x}_{t},t)-\nabla\log p_{0t}(\bm{x}_{t}|\bm{x}0)||_{2}^{2}\right],$ (3)
-where $\omega(t)$ is loss weights for the timestep $t$ and $p_{0t}(\bm{x}_{t}|\bm{x}_{0})$ is the transition density of $\bm{x}_{t}$ from the initial time to time $t$.
+where $\omega(t)$ is loss weights for $t$ and $p_{0t}(\bm{x}_{t}|\bm{x}_{0})$ is the transition density of $\bm{x}_{t}$ from the initial timestep 00 to $t$.
 This object can be interpreted as a noise-matching loss in DDPM [18], which predicts noise components in $\bm{x}_{t}$ and can be illustrated as $\int_{0}^{T}\mathbb{E}_{\bm{x}_{0}\sim p_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I})}[||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)-\epsilon||_{2}^{2}]\text{d}t$.
+This is regularly denoted as $\epsilon$-prediction parameterization [18, 20], and several other parameterizations including $F$-prediction [22, 26], score-prediction [57] and velocity-prediction [38] have been proposed.
 4 Observations
@@
 4.1 Analysis on the Task Difficulty in terms of Convergence
-Here, we analyze the convergence speed of loss and performance across timesteps $t$ by utilizing notable diffusion frameworks, DiT [20] and EDM [22].
-The details of experimental setups in our observation can be found in Appendix B.
-4.1.1 Convegence speed on loss.
+In this study, we analyze the convergence speed of loss and denoising performance across timesteps.
+To comprehensively cover various diffusion parameterizations, we utilized the notable frameworks DiT [20] for $\epsilon$-prediction, EDM [22] for $F$-prediction, SiT [38] for velocity prediction.
+Detailed descriptions of the experimental setups are provided in Appendix B.
+Convegence speed on loss.
 First, we analyze convergence characteristics of training loss across timesteps $t$.
-We divided whole timesteps $[0,T]$ into 20 uniformly divided intervals and trained 20 models $\{\mathrm{M}_{i}\}_{i=1}^{20}$ where $i$-th model learns denoising tasks in $[\frac{i-1}{20}T,\frac{i}{20}T]$.
-During training, we tracked the loss values through iterations and plotted their convergence speed by normalizing their value in Fig.
-0(a) and 0(c).
-As shown in the results, it is apparent that as $i$ increases towards $i=20$, the convergence accelerates in both DiT and EDM, suggesting that models learning larger timesteps can reach convergence more swiftly and reinforcing the notion that denoising tasks with larger timesteps are less difficult.
-4.1.2 Convegence speed on denoising performance.
-Although there is a known correlation between convergence speed of the loss function and task difficulties as indicated in prior curriculum learning literatures [26, 60], we delve deeper into the denoising performance convergence of 20 distinct models $\{\mathrm{M}_{i}\}_{i=1}^{20}$.
-To evaluate the performance of denoising tasks of each model, we generated samples where $\mathrm{M}_{i}$ was specifically employed for denoising within the timesteps that it was trained on, while the diffusion model learned whole timesteps handled denoising for the remaining timesteps.
-Then, the performance of the denoising capability of $\mathrm{M}_{i}$ can be quantitatively measured using the FID score [16], enabling us to observe the performance convergence of each model on its denoising tasks throughout the training process.
-Figures 0(b) and 0(d) depict this convergence.
-They illustrate that, similar to the loss convergence experiments, denoising performance converges more swiftly for models $\mathrm{M}_{i}$ with larger $i$ values, as observed in both DiT and EDM frameworks.
+We divided whole timesteps $[0,T]$ into 20 uniformly divided intervals and trained 20 models $\{\mathrm{M}_{i}\}_{i=1}^{20}$ where $i$-th model learns denoising tasks in $[\frac{i-1}{20}T,\frac{i}{20}T]$ for DiT and SiT, $[\mathrm{\Phi}^{-1}(\frac{i-1}{N}),\mathrm{\Phi}^{-1}(\frac{i}{N})]$ for EDM where $\mathrm{\Phi}^{-1}$ is the inverse cumulative distribution function of the Gaussian distribution.
+During training, we tracked the loss values through iterations and plotted their convergence speed by normalizing their value in Figs.
+1(a)-1(c).
+As shown in the results, it is apparent that as $i$ increases towards $i=20$, the convergence accelerates in both DiT, EDM, and SiT, suggesting that models learning larger timesteps can reach convergence more swiftly and reinforcing the notion that denoising tasks with larger timesteps are less difficult.
+Convegence speed on denoising performance.
+We also delve deeper into a convergence of denoising performance according to timesteps with 20 distinct models $\{\mathrm{M}_{i}\}_{i=1}^{20}$.
+To evaluate the performance of denoising tasks of each model, we generated samples where $\mathrm{M}_{i}$ was employed for denoising within the timesteps that it was trained on, while a diffusion model learned whole timesteps handled denoising for the remaining timesteps as in [11].
+Then, the performance of the denoising capability of $\mathrm{M}_{i}$ can be quantitatively measured using the FID score [16], enabling us to observe the performance convergence of each model on denoising tasks throughout the training process.
+Figures 1(d)-1(f) depict this convergence.
+They illustrate that, similar to loss convergence experiments, denoising performance converges more swiftly for models $\mathrm{M}_{i}$ with larger $i$ values, as observed across the DiT, EDM, and SiT.
 These results also suggest that models trained on later timesteps, indicated by larger $i$ values, achieve faster convergence, highlighting easier task difficulty at larger timesteps.
-(a) DiT (Loss)
-(b) DiT (FID)
-(c) EDM (Loss)
-(d) EDM (FID)
-Figure 1: Loss and FID scores plotted during training for each diffusion model $\mathrm{M}_{i}$ in DiT and EDM.
+(a) DiT (Loss convergence)
+(b) EDM (Loss convergence)
+(c) SiT (Loss convergence)
+(d) DiT (Task convergence)
+(e) EDM (Task convergence)
+(f) SiT (Task convergence)
+Figure 1: Loss and FID convergence plotted during training for each diffusion model $\mathrm{M}_{i}$ in DiT, EDM, and SiT.
 Since the loss scale for each model is different, we show the normalized value.
@@
 4.2 Exploration on Difficulties of Denoising Tasks
-Beyond empirical convergence metrics, we also delve into analyzing the relative entropy between $p_{t}$ and $p_{t-1}$ to gain a better understanding of task difficulties from a distributional viewpoint.
-The training of the diffusion model implicitly involves learning the distribution of the reverse process of the corresponding SDE shown in Eq.
-(1).
-To be specific, the transition probability of the reverse process is expressed as a conditional normal distribution whose mean parameter is modeled by neural networks, and the neural networks are thereby trained to learn the dynamics of the reverse process [18].
-Furthermore, the unconditional distribution of $\bm{x}_{t}$ can be obtained by marginalizing the transition densities over the prior distribution, indicating that information on the dynamics of the marginal distribution is fed to the neural networks [55].
+Beyond empirical convergence metrics, we also delve into analyzing the relative entropy between $p_{t}$ and $p_{t-1}$ to better understand task difficulties from a distributional perspective.
+The training of diffusion models implicitly involves learning the distribution of the reverse process of the corresponding SDE.
+To be specific, the transition probability of the reverse process is expressed as a conditional normal distribution whose mean parameter is modeled by neural networks, and they are thereby trained to learn the dynamics of the reverse process [18].
+Furthermore, an unconditional distribution of $\bm{x}_{t}$ can be obtained by marginalizing transition densities over the prior distribution, indicating that information on the dynamics of the marginal distribution is fed to neural networks [57].
+Figure 2: The KLD of $p_{t-1}$ from $p_{t}$ against denoising timestep. As the timestep increases, the dynamics decrease.
 To analyze the relationship between the dynamics of the unconditional distribution and the rate of loss convergence, we use the Kullback-Leibler (KL) divergence of $p_{t-1}$ to $p_{t}$, $D_{KL}(p_{t-1}||p_{t})$, as a quantitative measure.
 It is a pertinent divergence in that the training mechanism of diffusion models involves maximizing the likelihood of the reverse process.
-The KL divergence $D_{KL}(p_{t-1}||p_{t})$ is given by
-$D_{KL}\left(p_{t-1}||p_{t}\right)=\mathbb{E}_{\bm{x}\sim p_{t-1}}\left[\log\left(\frac{p_{t-1}(\bm{x})}{p_{t}(\bm{x})}\right)\right].$ (5)
-Moreover, the distribution $p_{t}$ of $\bm{x}$ at $t$ is expressed by
-$p_{t}(\bm{x}_{t})=\int p_{0t}(\bm{x}_{t}|\bm{x}_{0}=\bm{y})p_{0}(\bm{y})\text{d}y=\mathbb{E}_{\bm{x}_{0}\sim p_{0}}\left[p_{0t}\left(\bm{x}_{t}|\bm{x}_{0}\right)\right].$ (6)
-However, since the explicit density form of $p_{0}$ is unknown and it is computationally infeasible to estimate high-dimensional integrals in the Eqs.
-(5) and (6), we approximate them through unbiased estimators:
-$\hat{D}_{KL}\left(p_{t-1}||p_{t}\right)=\frac{1}{M}\sum_{\begin{subarray}{c}i\in\{1,2,\cdots,M\}\\ \bm{x}_{i}\sim p_{t-1}\end{subarray}}\log\left(\frac{p_{t-1}(\bm{x}_{i})}{p_{t}(\bm{x}_{i})}\right),$ (7)
-$\hat{p}_{t}(\bm{x}_{t})=\frac{1}{L}\sum_{\begin{subarray}{c}j\in\{1,2,\cdots,L\}\\ \bm{y}_{j}\sim p_{0}\end{subarray}}p_{0t}(\bm{x}_{t}|\bm{x}_{0}=\bm{y}_{j}),$ (8)
-where $\hat{D}_{KL}$ and $\hat{p}_{t}$ are unbiased estimators of $D_{KL}$ and $p_{t}$, respectively, and we choose Monte-Carlo estimators for them [10, 38].
+The KL divergence $D_{KL}(p_{t-1}||p_{t})$ is given by $D_{KL}\left(p_{t-1}||p_{t}\right)=\mathbb{E}_{\bm{x}\sim p_{t-1}}\left[\log\left(\frac{p_{t-1}(\bm{x})}{p_{t}(\bm{x})}\right)\right]$.
+Moreover, the distribution $p_{t}$ of $\bm{x}$ at $t$ is expressed by $p_{t}(\bm{x}_{t})=\int p_{0t}(\bm{x}_{t}|\bm{x}_{0}=\bm{y})p_{0}(\bm{y})\text{d}y=\mathbb{E}_{\bm{x}_{0}\sim p_{0}}\left[p_{0t}\left(\bm{x}_{t}|\bm{x}_{0}\right)\right]$.
+However, since the explicit density form of $p_{0}$ is unknown and it is computationally infeasible to estimate high-dimensional integrals, we approximate them through unbiased estimators (details in Appendix C).
 The empirical results of $D_{KL}(p_{t-1}||p_{t})$ for $64\times 64$ image data are given in Fig.
@@
 1.
-Figure 2: The KLD of $p_{t-1}$ from $p_{t}$ against denoising timestep. As the timestep increases, the dynamics decrease.
 This observation may stem from the inherent low-dimensional manifold of image data.
@@
 5.1 Design of Curriculum Learning in Diffusion Models
+Figure 3: The overview of our curriculum learning approach for diffusion models.
+(Left) We divide the timesteps into $N$ clusters, ${C_{1},...,C_{N}}$, with the difficulty of denoising tasks increasing from $C_{N}$ (easiest) to $C_{1}$ (hardest).
+(Right) As the curriculum progresses, learning accumulates harder task clusters, gradually increasing task difficulties.
 As we observed in Section 4, difficulties in denoising tasks increase as $t$ gets smaller.
-To utilize this in our training strategy for diffusion models, we adopt an easy-to-hard curriculum learning approach by partitioning the entire range of timesteps into $N$ clusters, denoted as $\{C_{i}\}_{i=1}^{N}$, where each cluster $C_{i}$ spans an interval $[l_{i},r_{i}]$, ensuring $l_{i}<r_{i}$ and continuity with $r_{i}=l_{i+1}$ for adjacent clusters.
-We then arrange for the denoising tasks to become progressively harder as we move from cluster $N$ to cluster $1$, ensuring $l_{i}<l_{i+1}$, with $l_{1}=0$ and $r_{N}=T$.
-During training, we introduce these clusters sequentially, starting from the least challenging (the $N$-th cluster) and advancing towards the most challenging (the $1$-st cluster), through $N$ distinct phases.
-Specifically, in the $n$-th stage, we jointly train the model on clusters $\{C_{N},C_{N-1},\ldots,C_{N-(n-1)}\}$.
-After this $N$-phase of curriculum learning, the model continues to learn across the entire range of timesteps, progressively converging towards a point of stability.
-The next consideration involves determining the boundaries for each cluster, $l_{i}$ and $r_{i}$.
+To utilize an easy-to-hard curriculum learning approach, we first divide the entire range of timesteps into $N$ clusters, denoted as $\{C_{i}\}_{i=1}^{N}$, where each cluster $C_{i}$ spans an interval $[l_{i},l_{i+1}]$, ensuring $l_{i}<l_{i+1}$, with $l_{1}=0$ and $l_{N+1}=T$, as shown on the left side of Fig.
+3.
+The curriculum for training is constructed by regarding these task clusters as unit tasks, starting from the least challenging (the $N$-th cluster $C_{N}$) and advancing towards the most difficult (the first cluster $C_{1}$), through $N$ distinct stages.
+Specifically, in the $n$-th curriculum stage, we jointly train the model with denoising tasks sampled from the clusters $\bigcup_{j=N-(n-1)}^{N}C_{j}$ as illustrated in the right side of Fig.
+3.
+The transition of the curriculum stages is determined by the pacing function, which will be discussed in the next section.
+After completing these $N$-stages of curriculum learning, the model continues to learn across the entire range of timesteps, $\bigcup_{j=1}^{N}C_{j}$, same as standard diffusion training.
+The next consideration involves determining the boundaries for each cluster $l_{i}$.
 A straightforward approach is to uniformly divide the entire timestep interval $[0,T]$ as $C_{i}=[\frac{(i-1)\cdot T}{N},\frac{i\cdot T}{N}]$ for $i=1,2,\ldots,N$.
@@
 Therefore, to address this issue more effectively, we adopt an SNR-based interval clustering technique as used in [11], which aligns the clustering with the actual changes in noise levels, potentially enhancing curriculum learning adaptability to varying noise conditions.
-In contrast, for EDM [23] which operates based on the noise level $\sigma$ rather than the timestep $t$, and where $\sigma$ is sampled from a log-normal distribution such that $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ during training, our clustering strategy for timesteps cannot be directly transposed.
+For EDM [24] which operates based on the noise level $\sigma$ rather than the timestep $t$, and where $\sigma$ is sampled from a log-normal distribution such that $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ during training, our clustering strategy for timesteps cannot be directly transposed.
 Given the log-normal distribution of $\sigma$, dividing it directly is impractical because $\sigma$ can extend over a wide range of values.
@@
 This method segments the noise levels into intervals by reflecting their probabilistic distribution.
-Algorithm 1 Curriculum Learning Input: Curriculum $\{C_{i}\}_{i=1}^{N}$, Pacing function $g$, Maximum patience $\tau_{\text{max}}$, Loss function $f$, Curriculum index $I_{\text{cur}}=N$, Best loss $L_{\text{best}}=\infty$, Model $M_{\theta}$ while $I_{\text{cur}}>0$ do     # Mini-batch sampling     $X\sim C_{I_{\text{cur}}}$     # Calculate Loss     $L_{\text{cur}}=f(M_{\theta}(X))$     # Update model     $\theta$ = $\theta$ $\nabla_{\theta}{L_{\text{cur}}}$     # Pacing function     $\tau_{\text{cur}},I_{\text{next}}=g(L_{\text{cur}},L_{\text{best}},\tau_{\text{cur}},\tau_{\text{max}},I_{\text{cur}})$     # Update curriculum     if $I_{\text{cur}}\neq I_{\text{next}}$ then        $I_{\text{cur}}=I_{\text{next}}$        $L_{\text{best}}=\infty$    # Update best loss     else if $L_{\text{cur}}<L_{\text{best}}$ then        $L_{\text{best}}=L_{\text{cur}}$     end if end while
-Algorithm 2 Pacing Function Input: Current loss $L_{\text{cur}}$, Best loss $L_{\text{best}}$, Current patience $\tau_{\text{cur}}$, Maximum patience $\tau_{\text{max}}$, Current curriculum index $I_{\text{cur}}$ Output: Updated patience, Updated curriculum index # Reset patience if $L_{\text{cur}}<L_{\text{best}}$ then     return 0, $I_{\text{cur}}$ else     # Proceed to next curriculum     if $\tau_{\text{cur}}+1>\tau_{\text{max}}$ then        return 0, $I_{\text{cur}}-1$    # Increase patience     else        return $p_{\text{cur}}+1,I_{\text{cur}}$     end if end if
 5.2 Pacing Strategy of Curriculum
-To effectively train the diffusion model according to the provided curriculum design, it is crucial to define a suitable pacing function for determining the transition of each $N$ distinct curriculum stage.
-Training for a fixed number of iterations for each curriculum is the simplest implementation (We also contain this method in experiments as ‘NaiveCL’ in Section 6.2).
-However, the convergence rate of each curriculum varies significantly, as demonstrated in Fig.
+To effectively train the diffusion model according to the provided curriculum design, it is crucial to define a suitable pacing function for determining the transition of each $N$ distinct curriculum.
+Training for a fixed number of iterations for each curriculum stage is the simplest implementation (We also contain this method in experiments as ‘NaiveCL’ in Section 6.2).
+However, the convergence rate of each curriculum phase varies significantly, as demonstrated in Fig.
 1.
@@
 [13].
-The proposed pacing function considers the loss of a mini-batch and maximum patience.
-If the loss does not improve within the maximum patience iterations, the current curriculum is terminated, and the subsequent curriculum is initiated.
-Here, the maximum patience is a fixed hyper-parameter and the detailed process is outlined in Algorithm 2.
-The overall curriculum learning procedure is presented in Algorithm 1.
+Our pacing function utilizes the training loss to determine transition moments and transitions to the next stage occur when the training loss converges at the current stage.
+Specifically, we introduce the maximum patience iteration $\tau$, and if the loss does not improve consecutively for $\tau$, the current curriculum stage is terminated, and the subsequent curriculum stage is initiated.
+Here, the maximum patience is a fixed hyper-parameter, and the detailed process and overall curriculum learning procedure are outlined in Algorithm 1 and 2 in Appendix D, respectively.
 6 Experimental Results
@@
 6.1 Experimental Setup
-Here, we provide experimental setups concisely due to space constraints. Detailed setups are presented in Appendix C.
-6.1.1 Evaluation protocols.
-For our comprehensive evaluation of various methods, we employed three distinct image-generation tasks: 1) Unconditional generation: We utilized the FFHQ dataset [23], which comprises 70,000 training images of human faces.
-2) Class-conditional generation: For this task, we employed two datasets: CIFAR-10 [28], which includes 60,000 training images, and ImageNet [7], containing 1,281,167 training images.
-In this setup, we applied classifier-free guidance [19].
-3) Text-to-Image generation: We used the MS-COCO dataset [33] , which contains 82,783 training images and 40,504 validation images.
-Each image in this dataset is accompanied by five text captions, facilitating the text-to-image generation task.
-6.1.2 Evaluation metrics.
-To evaluate the performance of models, we utilized three metrics: FID [16], IS [49], and Precision/Recall [30].
-Specifically, we applied FID and IS to measure sample quality, while Precision is used to assess quality further and Recall was utilized to evaluate the diversity of the generated samples in ImageNet setup.
-In other datasets, we employed FID to evaluate sample quality.
-6.1.3 Target models.
-We employed two exemplary diffusion architectures for experiments: DiT [42], which integrates latent diffusion models [46] with Transformer architectures [58], and EDM [22], which focuses on pixel-level diffusion utilizing UNet-based architectures [47].
+Here, we provide experimental setups concisely. Detailed setups are presented in Appendix E.
+Evaluation protocols.
+For our comprehensive evaluation of various methods, we employed three distinct image-generation tasks: 1) Unconditional generation with the FFHQ dataset [24], 2) Class-conditional generation with CIFAR-10 [28] and ImageNet [7] datasets, and 3) Text-to-Image generation with MS-COCO dataset [33].
+In 2) and 3) setups, we applied classifier-free guidance [19].
+Target models.
+We employed three exemplary diffusion architectures for experiments: DiT [44], which integrates latent diffusion models [48] with Transformer architectures [60] parameterized as $\epsilon$-prediction, EDM [22], which focuses on pixel-level diffusion utilizing UNet-based architectures [49] parameterized as $F$-prediction, and SiT [38] for score- and velocity-prediction.
 For the text-to-image generation, we incorporated a CLIP text encoder [46] as described in DTR [43].
@@
 In this section, we assess the effectiveness of our curriculum-based training approach.
-For a thorough comparison, we examine three distinct training variants, with further details provided in Appendix C: To achieve this, we compare three variants of training: 1) Vanilla: This term refers to diffusion models trained using conventional methods without any curriculum learning strategies; 2) NaiveCL: In this variant, we incorporate a basic curriculum learning strategy, which simply repeats the same number of iterations for each stage across an $N$-stage process and does not employ SNR-based clustering; 3) Ours: This denotes our proposed curriculum approach, which is designed to enhance the training process of diffusion models by systematically structuring the learning stages.
-6.2.1 Quantitative evaluation.
-As shown in Table 1, our approach was validated on various architectures—DiT [42] and EDM [22]— and tasks including unconditional, class-conditional, and text-to-image generation.
-To further contextualize our method, we evaluated the performance of NaiveCL, which does not leverage the SNR-based clustering and pacing function proposed in our approach.
-Our observations indicate that NaiveCL fails to consistently achieve improved performance, as it lacks the capability to incorporate task difficulty.
-In contrast, our method explicitly integrates the difficulty of denoising tasks, as detailed in Section 4.
-Consequently, our approach consistently demonstrates significant performance enhancements across all metrics on four datasets: FFHQ [23], ImageNet [7], CIFAR-10 [28], and MS-COCO [33], demonstrating that our approach works well data- and model-agnostic.
-Latent-Level Diffusion Model.
+For a thorough comparison, we examine three distinct training variants, with further details provided in Appendix E: To achieve this, we compare three variants of training: 1) Vanilla: This term refers to diffusion models trained using conventional methods without any curriculum learning strategies; 2) NaiveCL: In this variant, we incorporate a basic curriculum learning strategy, which simply repeats the same number of iterations for each stage across an $N$-stage process and does not employ SNR-based clustering; 3) Ours: This denotes our proposed curriculum approach, which is designed to enhance the training process of diffusion models by systematically structuring the learning stages.
+Quantitative evaluation.
+We quantitatively validate the effectiveness of our methods with various architectures-DiT [44], EDM [22], and SiT [38]- and tasks including unconditional, class-conditional, and text-to-image generation.
+Table 1 shows the results, confirming two empirical observations: 1) NaiveCL fails to consistently achieve improved performance compared to Vanilla, and 2) our approach outperforms both NaiveCL and Vanilla.
+Regarding the first observation, NaiveCL shows inconsistent improvements due to its lack of robust adaptation on incorporating task difficulties in various task conditions.
+In contrast, our method demonstrates superior performance across all scenarios by improving the clustering and pacing of curriculums.
+Consequently, our approach consistently achieves significant performance enhancements across all metrics on four datasets: FFHQ [24], ImageNet [7], CIFAR-10 [28], and MS-COCO [33], illustrating its effectiveness regardless of data or model used.
+Table 1: We evaluated unconditional image generation on FFHQ with DiT-B, EDM, and SiT-B, class-conditional image generation on ImageNet and CIFAR10 with DiT-L and EDM, respectively, and text-conditional image generation on MS-COCO with DiT-B.
+Note that our curriculum learning for diffusion models improves substantial performance without any additional parameters.
+$\epsilon$-prediction
 Model FFHQ 256$\times$256 ImageNet 256$\times$256 COCO 256$\times$256
@@
 DiT + Ours 7.55 8.18 186.37 0.79 0.47 7.51
-Pixel-Level Diffusion Model.
+$F$-prediction
 Model FFHQ 64$\times$64 CIFAR10 32$\times$32
@@
 EDM + Ours 2.71 2.44
-Table 1: We evaluated unconditional image generation on FFHQ with DiT-B and EDM, class-conditional image generation on ImageNet and CIFAR10 with DiT-L and EDM, respectively, and text-conditional image generation on MS-COCO with DiT-B.
-Note that our curriculum learning for diffusion models improves substantial performance without any additional parameters.
-6.2.2 Qualitative evaluation.
-Due to space constraints, we illustrate a detailed collection of generated examples in Appendix D.
-In summary, our curriculum methodology demonstrates a notable enhancement in the quality of the images produced, especially when compared to naive training.
+Velocity-prediction
+Model FFHQ 256$\times$256
+[…]
+SiT (Vanilla) 7.44
+SiT + NaiveCL 7.69
+SiT + Ours 6.95
+Score-prediction
+Model FFHQ 256$\times$256
+[…]
+SiT (Vanilla) 9.64
+SiT + NaiveCL 9.77
+SiT + Ours 9.15
+Qualitative evaluation.
+Due to space constraints, we illustrate a detailed collection of generated examples in Appendix F.
+In summary, our curriculum methodology demonstrates a notable enhancement in the quality of the images produced, when compared to NaiveCL and Vanilla.
 6.3 Analysis
-To elucidate the effectiveness of our curriculum approach, we hereby present a series of analytical studies. All the analysis is conducted by using the DiT-B model on the ImageNet dataset.
-6.3.1 Effects of $N$ and $\tau$.
+To elucidate our curriculum approach’s effectiveness, we present a series of analytical studies. All the analysis is conducted by using the DiT-B model on the ImageNet dataset.
 Figure 4: Ablation study on $N$ and $\tau$. We use DiT-B on ImageNet 256$\times$256.
+Effects of $N$ and $\tau$.
 We examined the robustness of the proposed curriculum training with respect to hyper-parameters: the number of clusters $N$ and the maximum patience $\tau$.
 As shown in Fig.
-3  our method consistently outperforms the vanilla model, and the best result is observed at $N=20,\tau=200$.
+4, our method consistently outperforms the vanilla model, and the best result is observed at $N=20,\tau=200$.
 It shows that as $\tau$ increases, it may lead to overtraining due to excessive iterations for each task, whereas as $\tau$ decreases, curriculums may not be sufficiently trained.
@@
 Conversely, with a decrease in $N$, tasks that should be in distinct clusters are learned together, forming a coarser cluster, which also leads to suboptimal outcomes.
-We utilized the optimal hyper-parameters for all subsequent experiments.
-6.3.2 Effects of Curriculum Design.
+Overall, our method outperforms vanilla training across a range of hyperparameters, demonstrating the robustness of our approach.
+Effects of Curriculum Design.
+Table 2: Comparative results on various curriculum designs.
 Class-Conditional ImageNet 256$\times$256.
@@
 (c2) + curriculum + SNR 22.96 75.98 0.62 0.52
-Table 2: Comparative results on various curriculum designs.
 In our curriculum design, we initially partitioned the entire set of timesteps into $N$ clusters using SNR-based clustering, organizing the curriculum from easy to hard clusters.
-To thoroughly assess the impact of each component, we conducted the ablation study as shown in Table.
-2.
-Firstly, we investigated the effect of curriculum learning via comparison with an anti-curriculum approach [13], which progresses from hard to easy clusters.
-While both training methods that consider task order (b2, c2) appear to enhance performance compared to vanilla training (a), anti-curriculum training cannot consistently guarantee performance improvement.
-To specifically assess the impact of anti-curriculum learning, we evaluate the performance of anti-curriculum learning without SNR-based clustering (b1), resulting in a performance decrease compared to vanilla.
-This can be attributed to the deviation from the observed appropriate task sequence, with the improvement seen in (b2) attributed to SNR-based clustering.
+To thoroughly assess the impact of each component, we conducted the ablation study as shown in Table 2.
+Firstly, we investigated the effect of curriculum learning via comparison with an anti-curriculum approach [13], which progresses from hard to easy clusters, unlike conventional curriculum learning.
+While both training methods in (b2) appear to enhance performance compared to vanilla training (a), anti-curriculum training cannot consistently guarantee performance improvement concerning the curriculum design as shown in (b1).
 In contrast, the proposed curriculum learning method (c1, c2) consistently exhibited performance improvement even with the uniformly partitioned clusters.
-Besides, with the finding that utilizing SNR-clustering was more effective, the clustering with the actual changes in noise levels enhanced the curriculum learning adaptability.
-We believe that by embracing more advanced clustering methods, we could gain further performance improvements.
-6.3.3 Visualization of curriculum.
-Figure 4: We visualized the curriculum transition and the corresponding loss across iterations ($N=20$). To make the loss graph more easily readable, the y-axis was truncated to 1.0.
+Besides, with findings that utilizing SNR-clustering was more effective, clustering with the actual changes in noise levels enhanced the curriculum learning adaptability.
+Visualization of curriculum.
+To gain deeper insights into the functioning of our curriculum pacing, we plotted loss metrics against curriculum phases, as illustrated in Fig.
+6.
+During the curriculum training, tasks progressively transition from the easiest to the most challenging, with varying amounts of iterations for each task based on the pacing function.
+The training loss decreased during each curriculum phase but increased after curriculum changes via the pacing function due to the inclusion of a newly added task in the updated curriculum.
+Additionally, as $\tau$ increases, the curriculum phases change more gradually, highlighting the role of $\tau$ in controlling the pace of curriculum transitions.
+Analysis on convergence speed.
+As demonstrated in previous works [3, 13], the adoption of curriculum learning can lead to faster convergence in model performance.
+To illustrate the efficacy of our approach in this regard, we plotted the FID, IS, precision, and recall calculated over 10,000 samples across the training iterations, as depicted in Fig.
+6.
+We observed the models trained through the proposed curriculum learning method converge faster than vanilla models, regardless of evaluation metrics.
+Notably, our approach achieves these improvements without requiring additional parameters or training iterations, thereby significantly saving time and computational resources.
 Figure 5: We visualized the curriculum transition and the corresponding loss across iterations ($N=20$). To make the loss graph more easily readable, the y-axis was truncated to 1.0.
 Figure 6: The models trained using the proposed curriculum learning approach demonstrate faster convergence compared to vanilla models, irrespective of evaluation metrics.
-To gain deeper insights into the functioning of our curriculum pacing, we plotted the loss metrics against the curriculum phases, as illustrated in Fig.
-5.
-During the curriculum training, tasks progressively transition from the easiest to the most challenging, with varying amounts of iterations for each task based on the pacing function.
-The training loss decreased during each curriculum phase but increased after curriculum changes via the pacing function due to the inclusion of a newly added task in the updated curriculum.
-In addition, as $\tau$ increases, it is confirmed that the curriculum phase slowly changes.
-6.3.4 Analysis on convergence speed.
-As demonstrated in previous works [3, 13], the adoption of curriculum learning can lead to faster convergence in model performance.
-To illustrate the efficacy of our approach in this regard, we plotted the FID, IS, precision, and recall calculated over 10,000 samples across the training iterations, as depicted in Fig.
-5.
-We observed the models trained through the proposed curriculum learning method converge faster than vanilla models, regardless of evaluation metrics.
-Note that our approach does not need any additional parameters and training iterations.
-6.3.5 Robustness on various sizes of models
+Effectiveness on various sizes of models
 To verify the generalizability of our method across different model sizes, we evaluated the performance gains achieved using our curriculum learning approach on various scales of the DiT model: DiT-S (small), DiT-B (base), and DiT-L (large).
@@
 4 shows that the proposed curriculum learning for diffusion model improves the performance regardless of the model size.
-Moreover, it is notable that larger models exhibit a more substantial performance enhancement: DiT-S has improved by 8% in terms of FID, while DiT-B and DiT-L have shown improvements of 24% and 27%, respectively.
-These findings validate the efficacy of our curriculum approach across a diverse range of model sizes, underscoring its robust generalizability.
-6.3.6 Additional experimental results
-Due to limited space, we present additional experimental results in Appendix E.
-These results also support the effectiveness of our method, emphasizing the importance of curriculum approaches in diffusion training.
-6.4 Orthogonality of Our Curriculum Approach
+Moreover, it is notable that larger models exhibit a more substantial performance enhancement: DiT-S improved by 8% in terms of FID, while DiT-B and DiT-L showed improvements of 24% and 27%, respectively.
+These findings validate the efficacy of our curriculum approach across a diverse range of model sizes, underscoring its generalizability to various model parameters.
+Orthogonality of Our Curriculum Approach
 Lastly, we illustrate the seamless integration of our method with sophisticated training techniques such as DTR [43] and MinSNR [14].
@@
 Consequently, the curriculum approach proves to be compatible with previous promising methods such as loss weighting (MinSNR) and architectural enhancements (DTR), demonstrating our orthogonality with recent diffusion techniques.
+Additional experimental results
+Due to limited space, we present additional experimental results in Appendix G.
+These results also support the effectiveness of our method, emphasizing the importance of curriculum approaches in diffusion training.
 Table 3: Note that the curriculum learning achieves consistent improvements across the model sizes.
+Class-Conditional ImageNet 256$\times$256.
 […]
+DiT-S/2 43.30 33.63 0.42 0.54
+DiT-S/2 + Ours 39.66 36.57 0.44 0.54
+DiT-B/2 30.27 60.06 0.55 0.52
+DiT-B/2 + Ours 22.96 75.98 0.62 0.52
+DiT-L/2 11.18 146.95 0.75 0.47
+DiT-L/2 + Ours 8.18 186.37 0.79 0.47
+Table 4: Note that the curriculum learning is compatible with the previous works such as the loss weighting (MinSNR) and architecture (DTR) study which, specified the multi-task learning for diffusion model.
+Class-Conditional ImageNet 256$\times$256.
+DiT-B/2 DiT-B/2 + Ours
 […]
-Table 3: Note that the curriculum learning achieves consistent improvements across the model sizes.
-Table 4: Note that the curriculum learning is compatible with the previous works such as loss weighting (MinSNR) and architecture (DTR) study which, specified the multi-task learning for diffusion model.
+Vanilla 30.27 60.06 0.55 0.52 22.96 75.98 0.62 0.52
+MinSNR [14] 21.88 88.12 0.63 0.49 19.36 101.35 0.67 0.49
+DTR [43] 15.77 89.89 0.68 0.52 15.33 91.39 0.68 0.52
 7 Conclusion
@@
 We believe that these observations might help reorganize the conflicts of previous works regarding denoising task difficulties.
-Building upon these insights, we propose a curriculum learning framework for the diffusion model, comprising curriculum design and pacing strategies.
+Building upon these insights, we propose a curriculum learning framework for diffusion models, comprising curriculum design and pacing strategies.
 Our experimental results convincingly demonstrate the efficacy of our approach across diverse diffusion model designs, datasets, and tasks.
-From these results, we emphasize that considering an order of learning denoising tasks is also a potential direction to improve the training of diffusion models.
+From these results, we emphasize that considering an order of learning denoising tasks is also a potential direction to improve training of diffusion models.
 In future research, for further enhancements, more advanced curriculum learning strategies such as self-pacing can be elaborated.
-A. Extended Related Work
+Appendix
+Appendix A Extended Related Work
 In this section, we review works related to analyzing diffusion model behaviors in each timestep but not covered in detail in Section 2.1.
@@
 For more stable training targets, they utilize weighted conditional scores with a reference batch.
-B. Detailed Experimental Setups for Observation
-In Section 4, we examined the difficulty of denoising tasks in terms of convergence with various models $\{\mathrm{M}\}_{i=1}^{20}$, which are trained within specific timesteps $[\frac{i-1}{20}T,\frac{i}{20}T]$.
+Appendix B Detailed Experimental Setups for Observation
+In Section 4, we examined the difficulty of denoising tasks in terms of convergence with various models $\{\mathrm{M}\}_{i=1}^{20}$, which are trained within specific timesteps $[\frac{i-1}{20}T,\frac{i}{20}T]$ for DiT and SiT, and $[\mathrm{\Phi}^{-1}(\frac{i-1}{N}),\mathrm{\Phi}^{-1}(\frac{i}{N})]$ for EDM where $\mathrm{\Phi}^{-1}$ is the inverse cumulative distribution function of the Gaussian distribution.
 For the DiT architecture, we employed the DiT-B/2, whereas for EDM, we used the DDPM++ architecture.
@@
 This training was conducted until both loss and performance converged.
+Configuration of training optimizers and learning rates are the same as setups in Section 6.
 To evaluate the performance of diffusion models through the FID score of the generated images, performing a recursive denoising task from $T$ to zero is necessary, complicating the assessment using only $\mathrm{M}_{i}$.
@@
 At the same time, the diffusion model is responsible for the denoising tasks across the entire range of timesteps.
-For this evaluation, we sampled 10,000 images using a DDIM sampler over 250 steps for DiT, and an Euler solver over 40 steps for the other models.
+For this evaluation, we sampled 10K images using a DDPM sampler over 250 steps for DiT and SiT, and an Euler solver over 40 steps for the other models.
 Vanilla
@@
 Ours
-Fig. A: Qualitative comparison between vanilla, naive curriculum, and ours on FFHQ dataset.
+Figure A: Qualitative comparison between vanilla, naive curriculum, and ours on the FFHQ dataset.
 Vanilla
@@
 Ostrich Clumber Arctic wolf Puffer American eagle Hummingbird Ridgeback
-Fig. B: Qualitative comparison between vanilla, naive curriculum, and ours on ImageNet dataset.
-C. Experiment Details
+Figure B: Qualitative comparison between vanilla, naive curriculum, and ours on ImageNet dataset.
+Appendix C Approximation of KL Divergence of $p_{t-1}$ and $p_{t}$.
+Here, we supplement the approximation of KL Divergence of $p_{t-1}$ and $p_{t}$ omitted in Section 4.2.
+To explore the difficulties of denoising tasks from the distributional viewpoint, we analyze the KL divergence of $p_{t-1}$ and $p_{t}$, $D_{KL}(p_{t-1}||p_{t})$.
+However, due to the unknown explicit density form of $p_{0}$, it is approximated through unbiased estimators as follows:
+$\hat{D}_{KL}\left(p_{t-1}||p_{t}\right)=\frac{1}{M}\sum_{\begin{subarray}{c}i\in\{1,2,\cdots,M\}\\ \bm{x}_{i}\sim p_{t-1}\end{subarray}}\log\left(\frac{p_{t-1}(\bm{x}_{i})}{p_{t}(\bm{x}_{i})}\right),$ (4)
+$\hat{p}_{t}(\bm{x}_{t})=\frac{1}{L}\sum_{\begin{subarray}{c}j\in\{1,2,\cdots,L\}\\ \bm{y}_{j}\sim p_{0}\end{subarray}}p_{0t}(\bm{x}_{t}|\bm{x}_{0}=\bm{y}_{j}),$ (5)
+where $\hat{D}_{KL}$ and $\hat{p}_{t}$ are unbiased estimators of $D_{KL}$ and $p_{t}$, respectively, and we choose Monte-Carlo estimators for them [10, 39].
+Appendix D Algorithm
+Due to the limited space of the main manuscript, we hereby present the step-by-step process of our method to supplement the details of our approach.
+The pacing function, which determines the moments to transit between curriculum stages is described in Algorithm 1.
+By incorporating this pacing function, the detailed procedure of our proposed curriculum learning method for training diffusion is illustrated in Algorithm 2.
+Algorithm 1 Pacing Function
+Input: Current loss $L_{\text{cur}}$, Best loss $L_{\text{best}}$, Current patience $\tau_{\text{cur}}$, Maximum patience $\tau_{\text{max}}$, Current curriculum index $I_{\text{cur}}$
+Output: Updated patience, Updated curriculum index
+# Reset patience
+if $L_{\text{cur}}<L_{\text{best}}$ then
+    return 0, $I_{\text{cur}}$
+else
+    # Proceed to next curriculum
+    if $\tau_{\text{cur}}+1>\tau_{\text{max}}$ then
+       return 0, $I_{\text{cur}}-1$
+   # Increase patience
+    else
+       return $p_{\text{cur}}+1,I_{\text{cur}}$
+    end if
+end if
+Algorithm 2 Curriculum Learning
+Input: Curriculum $\{C_{i}\}_{i=1}^{N}$, Pacing function $g$, Maximum patience $\tau_{\text{max}}$, Loss function $f$, Curriculum index $I_{\text{cur}}=N$, Best loss $L_{\text{best}}=\infty$, Model $M_{\theta}$
+while $I_{\text{cur}}>0$ do
+    # Mini-batch sampling
+    $X\sim C_{I_{\text{cur}}}$
+    # Calculate Loss
+    $L_{\text{cur}}=f(M_{\theta}(X))$
+    # Update model
+    $\theta$ = $\theta$ $\nabla_{\theta}{L_{\text{cur}}}$
+    # Pacing function
+    $\tau_{\text{cur}},I_{\text{next}}=g(L_{\text{cur}},L_{\text{best}},\tau_{\text{cur}},\tau_{\text{max}},I_{\text{cur}})$
+    # Update curriculum
+    if $I_{\text{cur}}\neq I_{\text{next}}$ then
+       $I_{\text{cur}}=I_{\text{next}}$
+       $L_{\text{best}}=\infty$
+   # Update best loss
+    else if $L_{\text{cur}}<L_{\text{best}}$ then
+       $L_{\text{best}}=L_{\text{cur}}$
+    end if
+end while
+Appendix E Details on Experimental Setups
+Evaluation metrics.
+To evaluate the performance of models, we utilized three metrics: FID [16], IS [51], and Precision/Recall [30].
+Specifically, we applied FID and IS to measure sample quality, while Precision is used to assess quality further and Recall was utilized to evaluate the diversity of the generated samples in ImageNet setup.
+In other datasets, we employed FID to evaluate sample quality.
+Unless otherwise mentioned, we sampled 50K samples for evaluation.
+In tasks involving conditional generation, including class-conditional image generation (e.g.
+CIFAR-10, ImageNet) and text-to-image conversion (e.g.
+MS-COCO), we adapted the classifier-free guidance [19] with a guidance scale of 1.5.
+Training details.
 For training diffusion models, we utilized the AdamW optimizer [36] with a constant learning rate of 0.0001, and weight decay was not applied.
@@
 EDM was trained using fp16, while the other models were trained using fp32.
-Unless otherwise mentioned, we sampled 50K samples for evaluation.
-In tasks involving conditional generation, including class-conditional image generation (e.g.
-CIFAR-10, ImageNet) and text-to-image conversion (e.g.
-MS-COCO), we adapted the classifier-free guidance [19] with a guidance scale of 1.5.
+We used 8 A100 GPUs for all experiments.
 Vanilla
@@
 Ours
-“A man is on a path riding a horse.”
-“Tower of a brick building with a clock on the front.”
-“A dog leaps through the air as it catches a frisbee.”
-“A person in white and red snow suit skiing on a slope.”
-“A group of young people playing a game of soccer.”
-“A young boy is smiling and he has food around him on a table.”
-“a small boy with a hat is standing on a surf board.”
-Fig. C: Qualitative comparison between vanilla, naive curriculum, and ours on MS-COCO dataset.
-D. Qualitative Results
-7.0.1 Qualitative Evaluation on the FFHQ Dataset.
+“A man is on a path riding a horse.” “Tower of a brick building with a clock on the front.” “A dog leaps through the air as it catches a frisbee.” “A person in white and red snow suit skiing on a slope.” “A group of young people playing a game of soccer.” “A young boy is smiling and he has food around him on a table.” “a small boy with a hat is standing on a surf board.”
+Figure C: Qualitative comparison between vanilla, naive curriculum, and ours on MS-COCO dataset.
+Appendix F Qualitative Results
+In this section, we present qualitative comparisons between three methods: 1) Vanilla, 2) NaiveCL, and 3) Ours, across the FFHQ, ImageNet, and MS-COCO datasets.
+All methods are evaluated using DiT-B models, and the final trained models generate all samples shown in the results.
+As shown in the results in the following subsections, our approach can synthesize more accurate and realistic images compared to Vanilla and NaiveCL.
+F.1 Qualitative Evaluation on the FFHQ Dataset.
 Figure A presents a qualitative analysis of the performance in unconditional facial image synthesis among the vanilla, the naive curriculum approach, and our method.
 Our approach demonstrates superior ability in generating realistic images.
-7.0.2 Qualitative Analysis on the ImageNet Dataset.
+F.2 Qualitative Analysis on the ImageNet Dataset.
 In the conditional image synthesis, we exhibit the outcomes generated by the vanilla, the naive curriculum strategy, and our proposed method.
 Figure B clearly shows that our methodology surpasses the competing approaches in terms of quality.
-7.0.3 Qualitative Assessment on the MS-COCO Dataset.
+F.3 Qualitative Assessment on the MS-COCO Dataset.
 To further substantiate the effectiveness of our proposed technique, we conduct a qualitative comparison of the results in the text-to-image generation task among the vanilla, the naive curriculum method, and our own approach, as depicted in Fig.
 C.
-E. Further Experimental Results
-E.1. Convergence Speed across Model Size
-Fig. D: We observed an increase in convergence speed across various model sizes when the proposed curriculum learning approach was applied.
+Appendix G Further Experimental Results
+G.1 Convergence Speed across Model Size
+Figure D: We observed an increase in convergence speed across various model sizes when the proposed curriculum learning approach was applied.
 By leveraging the advantages of curriculum learning in diffusion training, our method offers faster convergence than vanilla training.
@@
 These results also support the effectiveness of our method.
-E.2. Robustness on Noise Schedule
+G.2 Robustness on Noise Schedule
+Table A: Ablation study on noise scheduler. Note that our approach improves performance consistently across each scheduler.
 Class-Conditional ImageNet 256$\times$256.
@@
 Ours 16.03 87.66 0.69 0.53
-Table A: Ablation study on noise scheduler. Note that our approach improves performance consistently across each scheduler.
-For more comprehensive ablation study, we also trained the diffusion model with different noise schedule.
+For a more comprehensive ablation study, we also trained the diffusion model with different noise schedules.
 In contrast to cosine scheduling, the $\beta_{t}$ is set by uniformly dividing the interval $[0.0001,0.02]$, and the $C_{i}$ are obtained corresponding to SNR on a linear schedule.
 As shown in Table.
-A, our approach improves the performance with cosine and linear noise schedulers both.
-F. Negative Social Impacts
+A, our approach improves the performance with cosine and linear noise schedulers.
+G.3 Qualitative Results from DiT-L/2 with 2M iterations
+In Figures E-I, we present images generated by DiT-L using our curriculum training method for 2M iterations. The results demonstrate that our method produces highly realistic images.
+Figure E: Uncurated 256$\times$256 DiT-L/2 samples.
+Classifier-free guidanzce scale = 2.0.
+Class label = “golden retriever" (207)
+Figure F: Uncurated 256$\times$256 DiT-L/2 samples.
+Classifier-free guidance scale = 2.0.
+Class label = “panda" (388)
+Figure G: Uncurated 256$\times$256 DiT-L/2 samples.
+Classifier-free guidance scale = 2.0.
+Class label = “cliff drop-off" (972)
+Figure H: Uncurated 256$\times$256 DiT-L/2 samples.
+Classifier-free guidance scale = 2.0.
+Class label = “lake shore" (975)
+Figure I: Uncurated 256$\times$256 DiT-L/2 samples.
+Classifier-free guidance scale = 2.0.
+Class label = “lion" (291)
+Appendix H Broader Impacts
 Generative models, such as diffusion models, have the potential to significantly impact society, particularly through DeepFake applications and the use of biased datasets.
@@
 This situation may result in the dissemination of incorrect information and the manipulation of public opinion.
+Appendix I Limitations
+In this work, we demonstrated the varying difficulties of denoising tasks through empirical results on various diffusion frameworks and proposed a curriculum learning approach that effectively enhances diffusion model training.
+While we have shown the robustness of our method’s hyperparameters in improving vanilla diffusion training, there is potential for further improvement.
+Specifically, curriculum learning methods that utilize smaller hyperparameters and adjust dynamically based on the model itself could yield better results.
+We acknowledge the validity of this direction and consider it a promising avenue for future work.

### v3

@@
 1 Introduction
-Diffusion-based generative models [18, 52, 57] have achieved significant advancements in the realm of generative tasks, demonstrating notable success across various fields such as image [8], video [17, 15], and 3D [62, 35] generation.
-Specifically, their exceptional adaptability and promising performance in diverse image generation contexts, such as unconditional [22, 40], class-conditional [8], and text-conditional scenarios [2, 47], demonstrate their significant impact.
+Diffusion-based generative models (Ho et al., 2020, Sohl-Dickstein et al., 2015, Song et al., 2021) have achieved significant advancements in the realm of generative tasks, demonstrating notable success across various fields such as image (Dhariwal & Nichol, 2021), video (Ho et al., 2022a, Harvey et al., 2022), and 3D (Woo et al., 2023, Liu et al., 2023b) generation.
+Specifically, their exceptional adaptability and promising performance in diverse image generation contexts, such as unconditional (Karras et al., 2022, Nichol & Dhariwal, 2021), class-conditional (Dhariwal & Nichol, 2021), and text-conditional scenarios (Balaji et al., 2022, Ramesh et al., 2022), demonstrate their significant impact.
 Such achievements have led to a growing interest in further deepening the analysis and enhancing diffusion models.
-Diffusion models [18, 57] are designed to reverse the corruption of the data through the learning process at different noise levels and over multiple timesteps.
+Diffusion models (Ho et al., 2020, Song et al., 2021) are designed to reverse the corruption of the data through the learning process at different noise levels and over multiple timesteps.
 Recent works have delved into the learning of diffusion models across various noise levels and timesteps, revealing different stages of diffusion models.
 For example, Choi et al.
-[5] observe that when a diffusion model performs a denoising task from large to small timestep, it first generates coarse features, then gradually generates perceptually rich content, and later refines the details.
-Similar observation is also identified in text-to-image diffusion models [2].
-Besides this aspect, various studies have further explored the learning of diffusion models across timesteps and noise levels, elucidating their transition from denoising to generative functionalities [6], modular attributes [67], frequency characteristics [66, 31], trajectories [41], affinity [11], and variations of targets [64].
+(Choi et al., 2022) observe that when a diffusion model performs a denoising task from large to small timestep, it first generates coarse features, then gradually generates perceptually rich content, and later refines the details.
+Similar observation is also identified in text-to-image diffusion models (Balaji et al., 2022).
+Besides this aspect, various studies have further explored the learning of diffusion models across timesteps and noise levels, elucidating their transition from denoising to generative functionalities (Deja et al., 2022), modular attributes (Yue et al., 2024), frequency characteristics (Yang et al., 2023b, Lee et al., 2023), trajectories (Pan et al., 2024), affinity (Go et al., 2023a), and variations of targets (Xu et al., 2023).
 These observations have not only deepened understanding of diffusion models but have also directly contributed to improvement in diffusion models.
-Specifically, these insights are incorporated into their method design in various works, including loss functions [14, 5, 64], architectures [31, 2], accelerated sampling [41], representations [67], and guidance [12].
+Specifically, these insights are incorporated into their method design in various works, including loss functions (Hang et al., 2023, Xu et al., 2023), architectures (Lee et al., 2023, Balaji et al., 2022), accelerated sampling (Pan et al., 2024), representations (Yue et al., 2024), and guidance (Go et al., 2023b).
 Given the tangible benefits already realized from such studies, further in-depth analysis of diffusion models across timesteps and noise levels is crucial for uncovering insights and achieving unprecedented advancements in their capabilities.
 In this paper, to enrich the current understanding across various timesteps and noise levels, we investigate under-explored areas within diffusion models focusing on the task difficulties of denoising.
-Regarding denoising task difficulties, previous works speculate that denoising tasks across timesteps have different difficulties [32, 2], yet a detailed exploration of these variances remains sparse.
-Moreover, there exists a notable discrepancy among studies, with works identifying larger timesteps as more difficult [18, 14], while others argue that smaller timesteps pose greater difficulties [22, 9, 25].
+Regarding denoising task difficulties, previous works speculate that denoising tasks across timesteps have different difficulties (Li et al., 2023, Balaji et al., 2022), yet a detailed exploration of these variances remains sparse.
+Moreover, there exists a notable discrepancy among studies, with works identifying larger timesteps as more difficult (Ho et al., 2020, Hang et al., 2023), while others argue that smaller timesteps pose greater difficulties (Karras et al., 2022, Dockhorn et al., 2021, Kim et al., 2022).
 The discrepancy in difficulty across timesteps not only impedes the accurate interpretation of previous studies but also hinders the development of sophisticated training methods that properly utilize the timestep-wise variation in difficulty.
@@
 By integrating these, we confirm that denoising tasks at earlier timesteps are more difficult, indicated by slower convergence speeds and greater changes in relative entropy.
-Furthermore, building on these observations, we integrate an easy-to-hard learning scheme, a concept well-established in the curriculum learning literature [13, 27, 4, 61, 45], into the training process of diffusion models.
+Furthermore, building on these observations, we integrate an easy-to-hard learning scheme, a concept well-established in the curriculum learning literature (Hacohen & Weinshall, 2019, Kong et al., 2021, Chang et al., 2021, Wang et al., 2020, Pentina et al., 2015), into the training process of diffusion models.
 Specifically, we organize timesteps or noise levels into clusters and train the diffusion models with ascending levels of difficulty, moving from clusters categorized by higher to lower timesteps.
-After this curriculum process, models simultaneously learn whole timesteps as standard diffusion training [18, 57, 19] to reach the convergence point.
+After this curriculum process, models simultaneously learn whole timesteps as standard diffusion training (Ho et al., 2020, Song et al., 2021, Ho & Salimans, 2022) to reach the convergence point.
 Unlike conventional approaches where diffusion models are trained simultaneously across all timesteps, our method distinguishes itself by incorporating a sequential, order-aware training regime, reflecting an intended progression from easier to harder denoising tasks.
@@
 3) Orthogonality with Existing Improvements: Our approach is inherently model-agnostic, ensuring broad applicability across various diffusion models.
-Additionally, it can be integrated with advanced diffusion training techniques, such as loss weighting [5, 14, 11, 23].
+Additionally, it can be integrated with advanced diffusion training techniques, such as loss weighting (Choi et al., 2022, Hang et al., 2023, Go et al., 2023a, Karras et al., 2023).
 Finally, we empirically validate the advantages of our method by conducting comprehensive experiments across a variety of image-generation tasks.
-These include unconditional generation, class-conditional generation, and text-to-image generation, utilizing datasets such as FFHQ [24], ImageNet [7], and MS-COCO [33].
-By integrating our curriculum learning strategy into architectures— DiT [44], EDM [22], and SiT [38]—we demonstrate the efficacy of our approach in enhancing performance, accelerating convergence speed, and maintaining compatibility with existing techniques.
+These include unconditional generation, class-conditional generation, and text-to-image generation, utilizing datasets such as FFHQ (Karras et al., 2019), ImageNet (Deng et al., 2009), and MS-COCO (Lin et al., 2014).
+By integrating our curriculum learning strategy into architectures— DiT (Peebles & Xie, 2022), EDM (Karras et al., 2022), and SiT (Ma et al., 2024)—we demonstrate the efficacy of our approach in enhancing performance, accelerating convergence speed, and maintaining compatibility with existing techniques.
 2 Related Works
 2.1 Diffusion Models
-Diffusion models [18, 52, 57] are a group of generative models that create samples by utilizing a learned denoising process to noise.
-Several works have focused on improving diffusion models in various aspects, including model architectures [43, 8, 42], sampling speed [53, 37, 34], training objectives [14, 5, 11, 26, 38].
+Diffusion models (Ho et al., 2020, Sohl-Dickstein et al., 2015, Song et al., 2021) are a group of generative models that create samples by utilizing a learned denoising process to noise.
+Several works have focused on improving diffusion models in various aspects, including model architectures (Park et al., 2024b, Dhariwal & Nichol, 2021, Park et al., 2024a), sampling speed (Song et al., 2020, Lu et al., 2022, Liu et al., 2023a), training objectives (Hang et al., 2023, Choi et al., 2022, Go et al., 2023a, Kingma & Gao, 2023, Ma et al., 2024).
 These endeavors often involve investigating what diffusion models learn by dividing its process, aiming to enhance the performance of diffusion models.
-P2 [5] under-weights training loss functions at the clean-up stage from their observation that diffusion models learn coarse, perceptual, and removing noises at large, medium, and small timesteps.
-Ediff-I [2] observes that earlier sampling parts rely on conditions for generation, whereas later parts ignore the conditions.
+P2 (Choi et al., 2022) under-weights training loss functions at the clean-up stage from their observation that diffusion models learn coarse, perceptual, and removing noises at large, medium, and small timesteps.
+Ediff-I (Balaji et al., 2022) observes that earlier sampling parts rely on conditions for generation, whereas later parts ignore the conditions.
 They employ multiple denoisers to address the diverse characteristics of tasks associated with different parts of the sampling process.
-Moreover, various works have also investigated these aspects related to timesteps [6, 67, 66, 31, 41, 11, 64] (detailed illustrations can be found in Appendix A).
+Moreover, various works have also investigated these aspects related to timesteps (Deja et al., 2022, Yue et al., 2024, Yang et al., 2023b, Lee et al., 2023, Pan et al., 2024, Go et al., 2023a, Xu et al., 2023) (detailed illustrations can be found in Appendix A).
 While our study aligns with the above works, we analyze the under-explored aspect of denoising task difficulty.
@@
 Difficulties in denoising tasks in diffusion have been referred to by various works, but this aspect is not deeply explored.
-Several studies hypothesize that denoising tasks in diffusion encompass diverse difficulties [32, 2], and there have been conflicts regarding these difficulties between previous works.
-Certain studies consider denoising at larger noise levels and timesteps to be more difficult [18, 14], the focus is on the challenges associated with reconstructing data from substantial noise.
+Several studies hypothesize that denoising tasks in diffusion encompass diverse difficulties (Li et al., 2023, Balaji et al., 2022), and there have been conflicts regarding these difficulties between previous works.
+Certain studies consider denoising at larger noise levels and timesteps to be more difficult (Ho et al., 2020, Hang et al., 2023), the focus is on the challenges associated with reconstructing data from substantial noise.
 For instance, Hang et al.
-[14] articulate that while smaller timesteps (approaching zero) may require straightforward reconstructions, such strategies become less effective at higher noise levels or in larger timesteps.
+(Hang et al., 2023) articulate that while smaller timesteps (approaching zero) may require straightforward reconstructions, such strategies become less effective at higher noise levels or in larger timesteps.
 Similarly, Ho et al.
-[18] elucidate that their approach de-emphasizes loss terms at smaller timesteps to prioritize learning on the more challenging tasks at larger timesteps, thereby enhancing sample quality.
+(Ho et al., 2020) elucidate that their approach de-emphasizes loss terms at smaller timesteps to prioritize learning on the more challenging tasks at larger timesteps, thereby enhancing sample quality.
 Conversely, other studies argue that earlier timesteps or lower noise levels also present significant challenges.
 Karras et al.
-[22] suggest that detecting noise at low levels is challenging due to its minimal presence.
+Karras et al.
+(2022) suggest that detecting noise at low levels is challenging due to its minimal presence.
 Also, Kim et al.
-[25] illustrate the increasing difficulty and high variance in score estimation as timesteps approach zero, disturbing stable training of models.
+(Kim et al., 2022) illustrate the increasing difficulty and high variance in score estimation as timesteps approach zero, disturbing stable training of models.
 In line with these observations, Dockhorn et al.
-[9] build upon insights of [25], acknowledging the complexities at near zero timesteps, where score becomes highly complex and potentially unbounded.
+(Dockhorn et al., 2021) build upon insights of (Kim et al., 2022), acknowledging the complexities at near zero timesteps, where score becomes highly complex and potentially unbounded.
 In this work, we aim to resolve this conflict through an in-depth analysis of convergence properties and changes in relative entropy between consecutive probability distributions across timesteps.
 2.3 Curriculum Learning
-Curriculum learning [3, 13, 27], inspired by human learning patterns, is a method of training models in a structured order, starting with easier tasks [45] or examples [3] and gradually increasing difficulty.
-As pointed out by [3], curriculum learning formulation can be viewed as a continuation method [1], which starts from a smoother objective and gradually transformed into a less smooth version until it reaches the original objective function.
-Through this foundation, various works have achieved improved performance and faster convergence compared to standard training based on random mini-batches sampled uniformly [13, 27, 4, 61].
+Curriculum learning (Bengio et al., 2009, Hacohen & Weinshall, 2019, Kong et al., 2021), inspired by human learning patterns, is a method of training models in a structured order, starting with easier tasks (Pentina et al., 2015) or examples (Bengio et al., 2009) and gradually increasing difficulty.
+As pointed out by (Bengio et al., 2009), curriculum learning formulation can be viewed as a continuation method (Allgower & Georg, 2003), which starts from a smoother objective and gradually transformed into a less smooth version until it reaches the original objective function.
+Through this foundation, various works have achieved improved performance and faster convergence compared to standard training based on random mini-batches sampled uniformly (Hacohen & Weinshall, 2019, Kong et al., 2021, Chang et al., 2021, Wang et al., 2020).
 Curriculum learning primarily comprises two components: a curriculum scoring function, measuring the difficulty of tasks or examples, and a pacing function, modulating the speed of the curriculum progress.
-Regarding a curriculum score function, early studies have utilized human intuition for measuring difficulty, such as the complexity of geometric shapes in images [3] or the length of sequences [58].
-Recently, various works employ models to measure difficulty, including confidence of pre-trained models [13] and the loss of the current models [27].
-For the pacing function, a predefined pacing function has been employed, which involves training using a predetermined curriculum progression [13, 63].
-There are various forms of this and they can be generally represented as a function of training iteration [13, 63].
-Contrary to this, there have been proposals for pacing techniques dynamically adjusting based on the loss or performance of the current model during training [29, 21].
-In the diffusion model literature, curriculum learning has been utilized to organize the order of training data types based on prior knowledge of targeted generation tasks [59, 65].
+Regarding a curriculum score function, early studies have utilized human intuition for measuring difficulty, such as the complexity of geometric shapes in images (Bengio et al., 2009) or the length of sequences (Spitkovsky et al., 2010).
+Recently, various works employ models to measure difficulty, including confidence of pre-trained models (Hacohen & Weinshall, 2019) and the loss of the current models (Kong et al., 2021).
+For the pacing function, a predefined pacing function has been employed, which involves training using a predetermined curriculum progression (Hacohen & Weinshall, 2019, Wu et al., 2020).
+There are various forms of this and they can be generally represented as a function of training iteration (Hacohen & Weinshall, 2019, Wu et al., 2020).
+Contrary to this, there have been proposals for pacing techniques dynamically adjusting based on the loss or performance of the current model during training (Kumar et al., 2010, Jiang et al., 2014).
+In the diffusion model literature, curriculum learning has been utilized to organize the order of training data types based on prior knowledge of targeted generation tasks (Tang et al., 2023, Yang et al., 2023a).
 Tang et al.
-[59] sequentially train video diffusion models with lower resolution and FPS datasets before progressing to higher resolution and FPS datasets.
+(Tang et al., 2023) sequentially train video diffusion models with lower resolution and FPS datasets before progressing to higher resolution and FPS datasets.
 Similarly, Yang et al.
-[65] order text-to-sound generation data based on the number of events in audio clips, training diffusion models from lower to higher events datasets.
+(Yang et al., 2023a) order text-to-sound generation data based on the number of events in audio clips, training diffusion models from lower to higher events datasets.
 In contrast, our method explores the nature of denoising task difficulty in diffusion models and proposes a curriculum learning approach that progresses from easy to hard timesteps, deviating from the standard simultaneous training of all timesteps.
-Also, while consistency models [55, 54] adopt a curriculum approach to discretizing noise levels, progressively increasing the discretization steps of noise levels during training, we have distinct by exploring which noise level should be learned first and investigating the difficulty of denoising at each noise level.
+Also, while consistency models (Song et al., 2023, Song & Dhariwal, 2023) adopt a curriculum approach to discretizing noise levels, progressively increasing the discretization steps of noise levels during training, we have distinct by exploring which noise level should be learned first and investigating the difficulty of denoising at each noise level.
 3 Preliminaries
-In this section, we provide the necessary background on diffusion models [18, 52, 57].
+In this section, we provide the necessary background on diffusion models (Ho et al., 2020, Sohl-Dickstein et al., 2015, Song et al., 2021).
 Let $\bm{x}_{0}\in\mathbb{R}^{d}$ be a sample from the data distribution $p_{0}(\bm{x})$.
 The forward process of diffusion models transforms data $\bm{x}_{0}$ to latent $\bm{x}_{t\in[0,T]}$ by iteratively adding Gaussian noise.
-This can be formulated as a stochastic differential equation (SDE) [57] as $\text{d}\bm{x}_{t}=f(t)\bm{x}_{t}\text{d}t+g(t)\text{d}\bm{w}_{t}$, where $f(t)$ and $g(t)$ are drift and diffusion coefficients, and $\bm{w}_{t}$ is the standard Wiener process.
+This can be formulated as a stochastic differential equation (SDE) (Song et al., 2021) as $\text{d}\bm{x}_{t}=f(t)\bm{x}_{t}\text{d}t+g(t)\text{d}\bm{w}_{t}$, where $f(t)$ and $g(t)$ are drift and diffusion coefficients, and $\bm{w}_{t}$ is the standard Wiener process.
 The Gaussian transition kernel of this SDE is formulated as:
@@
 where $\bar{\bm{w}}_{t}$ and $\text{d}\bar{t}$ denote the reverse-time Wiener process and the infinitesimal reverse-time, respectively, with the actual data score $\nabla\log p_{t}(\bm{x}_{t})$.
-In most cases, a neural network $\epsilon_{\theta}$ having parameter $\theta$ is utilized to approximate this score function by learning the denoising tasks for each timestep $t$ from score matching loss $\mathcal{L}$ [56]:
+In most cases, a neural network $\epsilon_{\theta}$ having parameter $\theta$ is utilized to approximate this score function by learning the denoising tasks for each timestep $t$ from score matching loss $\mathcal{L}$ (Song & Ermon, 2019):
 $\mathcal{L}=\frac{1}{2}\int^{T}_{0}\mathcal{L}_{t}\text{d}t,\quad\mathcal{L}_{t}=\omega(t)\mathbb{E}_{\bm{x}_{t}\sim p_{0t}(\bm{x}_{t}|\bm{x}_{0}),\bm{x_{0}}\sim p_{0}}\left[||\epsilon_{\theta}(\bm{x}_{t},t)-\nabla\log p_{0t}(\bm{x}_{t}|\bm{x}0)||_{2}^{2}\right],$ (3)
 where $\omega(t)$ is loss weights for $t$ and $p_{0t}(\bm{x}_{t}|\bm{x}_{0})$ is the transition density of $\bm{x}_{t}$ from the initial timestep 00 to $t$.
-This object can be interpreted as a noise-matching loss in DDPM [18], which predicts noise components in $\bm{x}_{t}$ and can be illustrated as $\int_{0}^{T}\mathbb{E}_{\bm{x}_{0}\sim p_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I})}[||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)-\epsilon||_{2}^{2}]\text{d}t$.
-This is regularly denoted as $\epsilon$-prediction parameterization [18, 20], and several other parameterizations including $F$-prediction [22, 26], score-prediction [57] and velocity-prediction [38] have been proposed.
+This object can be interpreted as a noise-matching loss in DDPM (Ho et al., 2020), which predicts noise components in $\bm{x}_{t}$ and can be illustrated as $\int_{0}^{T}\mathbb{E}_{\bm{x}_{0}\sim p_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I})}[||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)-\epsilon||_{2}^{2}]\text{d}t$.
+This is regularly denoted as $\epsilon$-prediction parameterization (Ho et al., 2020, Jabri et al., 2022), and several other parameterizations including $F$-prediction (Karras et al., 2022, Kingma & Gao, 2023), score-prediction (Song et al., 2021) and velocity-prediction (Ma et al., 2024) have been proposed.
 4 Observations
@@
 Upon integrating our findings, we establish a key conclusion: the learning difficulty for denoising tasks escalates as the timestep $t$ decreases.
-4.1 Analysis on the Task Difficulty in terms of Convergence
+4.1 Analysis on the Task Difficulty in terms of Convergence Speed
 In this study, we analyze the convergence speed of loss and denoising performance across timesteps.
-To comprehensively cover various diffusion parameterizations, we utilized the notable frameworks DiT [20] for $\epsilon$-prediction, EDM [22] for $F$-prediction, SiT [38] for velocity prediction.
+To comprehensively cover various diffusion parameterizations, we utilized the notable frameworks DiT (Jabri et al., 2022) for $\epsilon$-prediction, EDM (Karras et al., 2022) for $F$-prediction, SiT (Ma et al., 2024) for velocity prediction.
 Detailed descriptions of the experimental setups are provided in Appendix B.
@@
 We also delve deeper into a convergence of denoising performance according to timesteps with 20 distinct models $\{\mathrm{M}_{i}\}_{i=1}^{20}$.
-To evaluate the performance of denoising tasks of each model, we generated samples where $\mathrm{M}_{i}$ was employed for denoising within the timesteps that it was trained on, while a diffusion model learned whole timesteps handled denoising for the remaining timesteps as in [11].
-Then, the performance of the denoising capability of $\mathrm{M}_{i}$ can be quantitatively measured using the FID score [16], enabling us to observe the performance convergence of each model on denoising tasks throughout the training process.
+To evaluate the performance of denoising tasks of each model, we generated samples where $\mathrm{M}_{i}$ was employed for denoising within the timesteps that it was trained on, while a diffusion model learned whole timesteps handled denoising for the remaining timesteps as in (Go et al., 2023a).
+Then, the performance of the denoising capability of $\mathrm{M}_{i}$ can be quantitatively measured using the FID score (Heusel et al., 2017), enabling us to observe the performance convergence of each model on denoising tasks throughout the training process.
 Figures 1(d)-1(f) depict this convergence.
@@
 4.2 Exploration on Difficulties of Denoising Tasks
+Figure 2: The KLD of $p_{t-1}$ from $p_{t}$ against denoising timestep. As the timestep increases, the dynamics decrease.
 Beyond empirical convergence metrics, we also delve into analyzing the relative entropy between $p_{t}$ and $p_{t-1}$ to better understand task difficulties from a distributional perspective.
 The training of diffusion models implicitly involves learning the distribution of the reverse process of the corresponding SDE.
-To be specific, the transition probability of the reverse process is expressed as a conditional normal distribution whose mean parameter is modeled by neural networks, and they are thereby trained to learn the dynamics of the reverse process [18].
-Furthermore, an unconditional distribution of $\bm{x}_{t}$ can be obtained by marginalizing transition densities over the prior distribution, indicating that information on the dynamics of the marginal distribution is fed to neural networks [57].
-Figure 2: The KLD of $p_{t-1}$ from $p_{t}$ against denoising timestep. As the timestep increases, the dynamics decrease.
+To be specific, the transition probability of the reverse process is expressed as a conditional normal distribution whose mean parameter is modeled by neural networks, and they are thereby trained to learn the dynamics of the reverse process (Ho et al., 2020).
+Furthermore, an unconditional distribution of $\bm{x}_{t}$ can be obtained by marginalizing transition densities over the prior distribution, indicating that information on the dynamics of the marginal distribution is fed to neural networks (Song et al., 2021).
 To analyze the relationship between the dynamics of the unconditional distribution and the rate of loss convergence, we use the Kullback-Leibler (KL) divergence of $p_{t-1}$ to $p_{t}$, $D_{KL}(p_{t-1}||p_{t})$, as a quantitative measure.
@@
 This observation may stem from the inherent low-dimensional manifold of image data.
-As is well-known (e.g., [50]), the image data is distributed on a relatively low-dimensional manifold with a narrow support and a highly peaked multi-modal structure.
+As is well-known (e.g., (Ruderman, 1994)), the image data is distributed on a relatively low-dimensional manifold with a narrow support and a highly peaked multi-modal structure.
 On the other hand, as Gaussian noise is iteratively added, the distribution of $\bm{x}_{t}$ approaches the independent Gaussian distribution in the ambient space.
@@
 This nature of the unconditional distribution may cause the relative entropy from $p_{t}$ to $p_{t-1}$ to decrease with $t$, indicating that it is more difficult to accurately represent the dynamics of the reverse process at small $t$.
+More discussion is in Appendix C.
 5 Methodology
 In Section 4, we observe that denoising tasks at smaller $t$ are more difficult to learn by models.
-From these order of difficulties in denoising tasks, we propose the incorporation strategy of an easy-to-hard training scheme, that has demonstrated its effectiveness in curriculum literature [3, 13, 27], for improving diffusion models’ training.
+From these order of difficulties in denoising tasks, we propose the incorporation strategy of an easy-to-hard training scheme, that has demonstrated its effectiveness in curriculum literature (Bengio et al., 2009, Hacohen & Weinshall, 2019, Kong et al., 2021), for improving diffusion models’ training.
 5.1 Design of Curriculum Learning in Diffusion Models
@@
 However, this method does not account for variations in noise levels across different timesteps.
-Therefore, to address this issue more effectively, we adopt an SNR-based interval clustering technique as used in [11], which aligns the clustering with the actual changes in noise levels, potentially enhancing curriculum learning adaptability to varying noise conditions.
-For EDM [24] which operates based on the noise level $\sigma$ rather than the timestep $t$, and where $\sigma$ is sampled from a log-normal distribution such that $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ during training, our clustering strategy for timesteps cannot be directly transposed.
+Therefore, to address this issue more effectively, we adopt an SNR-based interval clustering technique as used in (Go et al., 2023a), which aligns the clustering with the actual changes in noise levels, potentially enhancing curriculum learning adaptability to varying noise conditions.
+For EDM (Karras et al., 2019) which operates based on the noise level $\sigma$ rather than the timestep $t$, and where $\sigma$ is sampled from a log-normal distribution such that $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ during training, our clustering strategy for timesteps cannot be directly transposed.
 Given the log-normal distribution of $\sigma$, dividing it directly is impractical because $\sigma$ can extend over a wide range of values.
@@
 Hence, we propose adopting an adaptive number of iterations for each curriculum, akin to the varied exponential pacing approach explored by Hacohen et al.
-[13].
+(Hacohen & Weinshall, 2019).
 Our pacing function utilizes the training loss to determine transition moments and transitions to the next stage occur when the training loss converges at the current stage.
@@
 Evaluation protocols.
-For our comprehensive evaluation of various methods, we employed three distinct image-generation tasks: 1) Unconditional generation with the FFHQ dataset [24], 2) Class-conditional generation with CIFAR-10 [28] and ImageNet [7] datasets, and 3) Text-to-Image generation with MS-COCO dataset [33].
-In 2) and 3) setups, we applied classifier-free guidance [19].
+For our comprehensive evaluation of various methods, we employed three distinct image-generation tasks: 1) Unconditional generation with the FFHQ dataset (Karras et al., 2019), 2) Class-conditional generation with CIFAR-10 (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009) datasets, and 3) Text-to-Image generation with MS-COCO dataset (Lin et al., 2014).
+In 2) and 3) setups, we applied classifier-free guidance (Ho & Salimans, 2022).
 Target models.
-We employed three exemplary diffusion architectures for experiments: DiT [44], which integrates latent diffusion models [48] with Transformer architectures [60] parameterized as $\epsilon$-prediction, EDM [22], which focuses on pixel-level diffusion utilizing UNet-based architectures [49] parameterized as $F$-prediction, and SiT [38] for score- and velocity-prediction.
-For the text-to-image generation, we incorporated a CLIP text encoder [46] as described in DTR [43].
+We employed three exemplary diffusion architectures for experiments: DiT (Peebles & Xie, 2022), which integrates latent diffusion models (Rombach et al., 2022) with Transformer architectures (Vaswani et al., 2017) parameterized as $\epsilon$-prediction, EDM (Karras et al., 2022), which focuses on pixel-level diffusion utilizing UNet-based architectures (Ronneberger et al., 2015) parameterized as $F$-prediction, and SiT (Ma et al., 2024) for score- and velocity-prediction.
+For the text-to-image generation, we incorporated a CLIP text encoder (Radford et al., 2021) as described in DTR (Park et al., 2024b).
 6.2 Comparative Results
@@
 Quantitative evaluation.
-We quantitatively validate the effectiveness of our methods with various architectures-DiT [44], EDM [22], and SiT [38]- and tasks including unconditional, class-conditional, and text-to-image generation.
-Table 1 shows the results, confirming two empirical observations: 1) NaiveCL fails to consistently achieve improved performance compared to Vanilla, and 2) our approach outperforms both NaiveCL and Vanilla.
-Regarding the first observation, NaiveCL shows inconsistent improvements due to its lack of robust adaptation on incorporating task difficulties in various task conditions.
-In contrast, our method demonstrates superior performance across all scenarios by improving the clustering and pacing of curriculums.
-Consequently, our approach consistently achieves significant performance enhancements across all metrics on four datasets: FFHQ [24], ImageNet [7], CIFAR-10 [28], and MS-COCO [33], illustrating its effectiveness regardless of data or model used.
 Table 1: We evaluated unconditional image generation on FFHQ with DiT-B, EDM, and SiT-B, class-conditional image generation on ImageNet and CIFAR10 with DiT-L and EDM, respectively, and text-conditional image generation on MS-COCO with DiT-B.
@@
 SiT + Ours 9.15
+Table 2: Evaluating the effectiveness of curriculum learning with extended training iterations on the ImageNet 256x256 dataset using the DiT-L architecture.
+[…]
+DiT (Vanilla) 11.18 7.84
+DiT + Ours 8.18 6.24
+We quantitatively validate the effectiveness of our methods with various architectures-DiT (Peebles & Xie, 2022), EDM (Karras et al., 2022), and SiT (Ma et al., 2024)- and tasks including unconditional, class-conditional, and text-to-image generation.
+Table 1 shows the results, confirming two empirical observations: 1) NaiveCL fails to consistently achieve improved performance compared to Vanilla, and 2) our approach outperforms both NaiveCL and Vanilla.
+Regarding the first observation, NaiveCL shows inconsistent improvements due to its lack of robust adaptation on incorporating task difficulties in various task conditions.
+In contrast, our method demonstrates superior performance across all scenarios by improving the clustering and pacing of curriculums.
+Consequently, our approach consistently achieves significant performance enhancements across all metrics on four datasets: FFHQ (Karras et al., 2019), ImageNet (Deng et al., 2009), CIFAR-10 (Krizhevsky et al., 2009), and MS-COCO (Lin et al., 2014), illustrating its effectiveness regardless of data or model used.
+Showing the results of longer training might demonstrate the robustness of our method in more extended training scenarios.
+We trained DiT-L/2 with 2M iterations and reported the results in Table 2.
+Our model consistently outperformed the baseline, demonstrating its effectiveness even in prolonged training.
+Therefore, our method proves to be robust and effective for longer training durations.
 Qualitative evaluation.
@@
 In summary, our curriculum methodology demonstrates a notable enhancement in the quality of the images produced, when compared to NaiveCL and Vanilla.
+Figure 4: Ablation study on $N$ and $\tau$. We use DiT-B on ImageNet 256$\times$256.
 6.3 Analysis
 To elucidate our curriculum approach’s effectiveness, we present a series of analytical studies. All the analysis is conducted by using the DiT-B model on the ImageNet dataset.
-Figure 4: Ablation study on $N$ and $\tau$. We use DiT-B on ImageNet 256$\times$256.
 Effects of $N$ and $\tau$.
@@
 To thoroughly assess the impact of each component, we conducted the ablation study as shown in Table 3.
-Firstly, we investigated the effect of curriculum learning via comparison with an anti-curriculum approach [13], which progresses from hard to easy clusters, unlike conventional curriculum learning.
+Firstly, we investigated the effect of curriculum learning via comparison with an anti-curriculum approach (Hacohen & Weinshall, 2019), which progresses from hard to easy clusters, unlike conventional curriculum learning.
 While both training methods in (b2) appear to enhance performance compared to vanilla training (a), anti-curriculum training cannot consistently guarantee performance improvement concerning the curriculum design as shown in (b1).
@@
 Analysis on convergence speed.
-As demonstrated in previous works [3, 13], the adoption of curriculum learning can lead to faster convergence in model performance.
+As demonstrated in previous works (Bengio et al., 2009, Hacohen & Weinshall, 2019), the adoption of curriculum learning can lead to faster convergence in model performance.
 To illustrate the efficacy of our approach in this regard, we plotted the FID, IS, precision, and recall calculated over 10,000 samples across the training iterations, as depicted in Fig.
@@
 Orthogonality of Our Curriculum Approach
-Lastly, we illustrate the seamless integration of our method with sophisticated training techniques such as DTR [43] and MinSNR [14].
+Lastly, we illustrate the seamless integration of our method with sophisticated training techniques such as DTR (Park et al., 2024b) and MinSNR (Hang et al., 2023).
 Initially, we observed that each sophisticated method yields a superior performance compared to the vanilla method.
 Meanwhile, as shown in Table.
-4, the performance is significantly enhanced when we apply the proposed curriculum learning.
+5, the performance is significantly enhanced when we apply the proposed curriculum learning.
 Consequently, the curriculum approach proves to be compatible with previous promising methods such as loss weighting (MinSNR) and architectural enhancements (DTR), demonstrating our orthogonality with recent diffusion techniques.
@@
 DiT-L/2 + Ours 8.18 186.37 0.79 0.47
+DiT-XL/2 9.40 166.83 0.77 0.49
+DiT-XL/2 + Ours 7.57 234.93 0.82 0.48
 Table 5: Note that the curriculum learning is compatible with the previous works such as the loss weighting (MinSNR) and architecture (DTR) study which, specified the multi-task learning for diffusion model.
@@
 Vanilla 30.27 60.06 0.55 0.52 22.96 75.98 0.62 0.52
-MinSNR [14] 21.88 88.12 0.63 0.49 19.36 101.35 0.67 0.49
-DTR [43] 15.77 89.89 0.68 0.52 15.33 91.39 0.68 0.52
+MinSNR (Hang et al., 2023) 21.88 88.12 0.63 0.49 19.36 101.35 0.67 0.49
+DTR (Park et al., 2024b) 15.77 89.89 0.68 0.52 15.33 91.39 0.68 0.52
 7 Conclusion
@@
 Appendix A Extended Related Work
+A.1 Analyzing Diffusion Model Behaviors in Each Timestep
 In this section, we review works related to analyzing diffusion model behaviors in each timestep but not covered in detail in Section 2.1.
 Deja et al.
-[6] explore denoising during the backward diffusion process and observe that transition from denoising to generation exists in the backward process.
+(Deja et al., 2022) explore denoising during the backward diffusion process and observe that transition from denoising to generation exists in the backward process.
 Go et al.
-[11] investigate the affinity between denoising tasks, showing that temporal proximal denoising tasks exhibit higher task affinity.
+(Go et al., 2023a) investigate the affinity between denoising tasks, showing that temporal proximal denoising tasks exhibit higher task affinity.
 Then, they also observe that simultaneously learning all denoising tasks by one model suffers from negative transfer.
@@
 Lee et al.
-[31] analyze frequency characteristics according to timesteps and observe that high-frequency components are lost as timesteps increase.
+(Lee et al., 2023) analyze frequency characteristics according to timesteps and observe that high-frequency components are lost as timesteps increase.
 From this observation, they propose a multi-architecture multi-experts diffusion model, which utilizes multiple denoiser models specialized in each timestep interval but utilizes a transformer-like model as the timestep increases.
 From observations that smaller and larger models produce similar latent noise, Pan et al.
-[41] propose T-Stitch, which leverages a pre-trained smaller model at the beginning of the backward process to accelerate the sampling speed.
+(Pan et al., 2024) propose T-Stitch, which leverages a pre-trained smaller model at the beginning of the backward process to accelerate the sampling speed.
 Xu et al.
-[64] investigate the average trace-of-covariance of training targets according to timesteps, showing that it peaks in the intermediate timesteps, causing unstable training targets.
+(Xu et al., 2023) investigate the average trace-of-covariance of training targets according to timesteps, showing that it peaks in the intermediate timesteps, causing unstable training targets.
 For more stable training targets, they utilize weighted conditional scores with a reference batch.
+A.2 Easy-to-hard training Strategy
+Progressive distillation (Salimans & Ho, 2022) focuses on reducing the number of sampling steps by training the model to progressively skip more steps, while cascaded diffusion (Ho et al., 2022b) aims to improve sample quality by progressively increasing the image resolution during training.
+Both methods concentrate on altering the model’s behavior or structure to tackle specific challenges, such as efficiency or resolution enhancement.
+In contrast, our work identifies trends in task difficulty across timestep-wise denoising tasks and leverages these findings to propose an easy-to-hard training scheme.
+This training strategy directly addresses the order and structure of the learning process, optimizing task sequencing to enhance performance.
+This distinction emphasizes that our approach is fundamentally different from these methods, as it addresses a unique aspect of diffusion model training.
 Appendix B Detailed Experimental Setups for Observation
@@
 For the DiT architecture, we employed the DiT-B/2, whereas for EDM, we used the DDPM++ architecture.
-Both DiT and EDM models were trained on the FFHQ dataset, with a batch size of 256, for approximately 20,000 and 4,000 iterations (equivalent to processing 1 million images), respectively.
+Both DiT and EDM models were trained on the FFHQ dataset, with a batch size of 256, for approximately 20,000 iterations and 4,000 kimg iterations (equivalent to processing 1 million images), respectively.
 This training was conducted until both loss and performance converged.
+As illustrated in Fig.
+A, we additionally plotted the iterations of each timestep interval when their loss values start to oscillate.
+We measured this by counting the number of times the loss value increased after the step reached 100.
+As shown in the results, losses of all timestep intervals are stabilized within 20K iterations, while the lower timesteps reach this point more slowly.
+This also suggests that the convergence speed of lower timesteps tends to exhibit a slower regime.
+To examine specifically at the observation of convergence, we also analyzed the convergence speed on the ImageNet dataset.
+As shown in Fig.
+B, we obtained similar results as on the FFHQ dataset.
 Configuration of training optimizers and learning rates are the same as setups in Section 6.
+Figure A: Converged points are plotted during training for each diffusion model $\mathrm{M}_{i}$ in SiT.
+Figure B: Loss convergence plotted during training for each diffusion model $\mathrm{M}_{i}$ in DiT on ImageNet dataset.
 To evaluate the performance of diffusion models through the FID score of the generated images, performing a recursive denoising task from $T$ to zero is necessary, complicating the assessment using only $\mathrm{M}_{i}$.
-Following [11], we generated samples where $\mathrm{M}_{i}$ was specifically utilized for denoising within its trained intervals.
+Following (Go et al., 2023a), we generated samples where $\mathrm{M}_{i}$ was specifically utilized for denoising within its trained intervals.
 At the same time, the diffusion model is responsible for the denoising tasks across the entire range of timesteps.
 For this evaluation, we sampled 10K images using a DDPM sampler over 250 steps for DiT and SiT, and an Euler solver over 40 steps for the other models.
-Vanilla
-Naive
-Ours
-Figure A: Qualitative comparison between vanilla, naive curriculum, and ours on the FFHQ dataset.
-Vanilla
-Naive
-Ours
-Ostrich Clumber Arctic wolf Puffer American eagle Hummingbird Ridgeback
-Figure B: Qualitative comparison between vanilla, naive curriculum, and ours on ImageNet dataset.
 Appendix C Approximation of KL Divergence of $p_{t-1}$ and $p_{t}$.
-Here, we supplement the approximation of KL Divergence of $p_{t-1}$ and $p_{t}$ omitted in Section 4.2.
-To explore the difficulties of denoising tasks from the distributional viewpoint, we analyze the KL divergence of $p_{t-1}$ and $p_{t}$, $D_{KL}(p_{t-1}||p_{t})$.
+Here, we supplement the approximation of KL Divergence of $p_{t-1}$ and $p_{t}$ omitted in Section 4.2.To explore the difficulties of denoising tasks from the distributional viewpoint, we analyze the KL divergence of $p_{t-1}$ and $p_{t}$, $D_{KL}(p_{t-1}||p_{t})$.
 However, due to the unknown explicit density form of $p_{0}$, it is approximated through unbiased estimators as follows:
@@
 $\hat{p}_{t}(\bm{x}_{t})=\frac{1}{L}\sum_{\begin{subarray}{c}j\in\{1,2,\cdots,L\}\\ \bm{y}_{j}\sim p_{0}\end{subarray}}p_{0t}(\bm{x}_{t}|\bm{x}_{0}=\bm{y}_{j}),$ (5)
-where $\hat{D}_{KL}$ and $\hat{p}_{t}$ are unbiased estimators of $D_{KL}$ and $p_{t}$, respectively, and we choose Monte-Carlo estimators for them [10, 39].
+where $\hat{D}_{KL}$ and $\hat{p}_{t}$ are unbiased estimators of $D_{KL}$ and $p_{t}$, respectively, and we choose Monte-Carlo estimators for them (Glasserman, 2004, McLeish, 2011).
+We sampled 5,000 images to approximate the KL divergence, which is enough for Monte-Carlo sampling and might be no changes for larger samples.
+Despite the large amount of samples, the exploding appearance observed in Fig.
+2 when $t$ is close to zero is due to the characteristics of the data distribution.
+The image data distribution has narrow support (roughly speaking, it is non-zero only within a narrow range) (Ruderman & Bialek, 1993, Karras et al., 2024).
+As $t$ increases, information about the original data distribution gradually diminishes with the modes in the distribution of $x_{t}$ vanishing towards zero.
+Given this, when $t$ is close to zero (i.e.
+when the distribution of $x_{t}$ is still analogous to the original data distribution), the narrow support and the tendency to move towards zero give rise to a region where $p_{t-1}$ does not overlap with $p_{t}$.
+Consequently, when calculating the KL divergence $D_{KL}(p_{t-1}||p_{t})=E_{x\sim p_{t-1}}[\log(\frac{p_{t-1}({x})}{p_{t}({x})})]$, $x_{t-1}$ potentially falls outside the support of $p_{t}$, which leads to $p_{t}(x_{t-1})=0$ and numerical unstability.
+On the other hand, as $t$ increases, the accumulated noise broadens the support of $x$’s distribution, reducing the occurrence of zero values and stabilizing the numerical estimation.
 Appendix D Algorithm
@@
 Evaluation metrics.
-To evaluate the performance of models, we utilized three metrics: FID [16], IS [51], and Precision/Recall [30].
+To evaluate the performance of models, we utilized three metrics: FID (Heusel et al., 2017), IS (Salimans et al., 2016), and Precision/Recall (Kynkäänniemi et al., 2019).
 Specifically, we applied FID and IS to measure sample quality, while Precision is used to assess quality further and Recall was utilized to evaluate the diversity of the generated samples in ImageNet setup.
@@
 CIFAR-10, ImageNet) and text-to-image conversion (e.g.
-MS-COCO), we adapted the classifier-free guidance [19] with a guidance scale of 1.5.
+MS-COCO), we adapted the classifier-free guidance (Ho & Salimans, 2022) with a guidance scale of 1.5.
 Training details.
-For training diffusion models, we utilized the AdamW optimizer [36] with a constant learning rate of 0.0001, and weight decay was not applied.
+For training diffusion models, we utilized the AdamW optimizer (Loshchilov & Hutter, 2017) with a constant learning rate of 0.0001, and weight decay was not applied.
 The exponential moving average (EMA) to the model’s weights was used to stabilize the training and the decay ratio was set to 0.9999.
 The batch size was set to 256, and we augmented the training data by a horizontal flip.
-While the diffusion timestep $T$ was configured as 1,000 for all experiments, we trained for 100K iterations for the FFHQ dataset [24], and 400K iterations for the ImageNet dataset [7] and MS-COCO dataset [33].
+While the diffusion timestep $T$ was configured as 1,000 for all experiments, we trained for 100K iterations for the FFHQ dataset (Karras et al., 2019), and 400K iterations for the ImageNet dataset (Deng et al., 2009) and MS-COCO dataset (Lin et al., 2014).
 The number of clusters $N$ was 20 unless otherwise specified.
@@
 Ours
+Figure C: Qualitative comparison between vanilla, naive curriculum, and ours on the FFHQ dataset.
+Vanilla
+Naive
+Ours
+ Ostrich Clumber Arctic wolf Puffer American eagle Hummingbird Ridgeback
+Figure D: Qualitative comparison between vanilla, naive curriculum, and ours on ImageNet dataset.
+Vanilla
+Naive
+Ours
 “A man is on a path riding a horse.” “Tower of a brick building with a clock on the front.” “A dog leaps through the air as it catches a frisbee.” “A person in white and red snow suit skiing on a slope.” “A group of young people playing a game of soccer.” “A young boy is smiling and he has food around him on a table.” “a small boy with a hat is standing on a surf board.”
-Figure C: Qualitative comparison between vanilla, naive curriculum, and ours on MS-COCO dataset.
+Figure E: Qualitative comparison between vanilla, naive curriculum, and ours on MS-COCO dataset.
 Appendix F Qualitative Results
@@
 F.1 Qualitative Evaluation on the FFHQ Dataset.
-Figure A presents a qualitative analysis of the performance in unconditional facial image synthesis among the vanilla, the naive curriculum approach, and our method.
+Figure C presents a qualitative analysis of the performance in unconditional facial image synthesis among the vanilla, the naive curriculum approach, and our method.
 Our approach demonstrates superior ability in generating realistic images.
@@
 In the conditional image synthesis, we exhibit the outcomes generated by the vanilla, the naive curriculum strategy, and our proposed method.
-Figure B clearly shows that our methodology surpasses the competing approaches in terms of quality.
+Figure D clearly shows that our methodology surpasses the competing approaches in terms of quality.
 F.3 Qualitative Assessment on the MS-COCO Dataset.
 To further substantiate the effectiveness of our proposed technique, we conduct a qualitative comparison of the results in the text-to-image generation task among the vanilla, the naive curriculum method, and our own approach, as depicted in Fig.
-C.
+E.
 Appendix G Further Experimental Results
 G.1 Convergence Speed across Model Size
-Figure D: We observed an increase in convergence speed across various model sizes when the proposed curriculum learning approach was applied.
+Figure F: We observed an increase in convergence speed across various model sizes when the proposed curriculum learning approach was applied.
 By leveraging the advantages of curriculum learning in diffusion training, our method offers faster convergence than vanilla training.
 To further investigate this aspect, we measured FID-10K through training iterations for DiT-S and DiT-L.
-Figure D describes the results, showing that our curriculum approach achieves faster convergence in both models.
+Figure F describes the results, showing that our curriculum approach achieves faster convergence in both models.
 These results also support the effectiveness of our method.
@@
 G.3 Qualitative Results from DiT-L/2 with 2M iterations
-In Figures E-I, we present images generated by DiT-L using our curriculum training method for 2M iterations. The results demonstrate that our method produces highly realistic images.
-Figure E: Uncurated 256$\times$256 DiT-L/2 samples.
+In Figures G-K, we present images generated by DiT-L using our curriculum training method for 2M iterations. The results demonstrate that our method produces highly realistic images.
+Figure G: Uncurated 256$\times$256 DiT-L/2 samples.
 Classifier-free guidanzce scale = 2.0.
 Class label = “golden retriever" (207)
-Figure F: Uncurated 256$\times$256 DiT-L/2 samples.
+Figure H: Uncurated 256$\times$256 DiT-L/2 samples.
 Classifier-free guidance scale = 2.0.
 Class label = “panda" (388)
-Figure G: Uncurated 256$\times$256 DiT-L/2 samples.
+Figure I: Uncurated 256$\times$256 DiT-L/2 samples.
 Classifier-free guidance scale = 2.0.
 Class label = “cliff drop-off" (972)
-Figure H: Uncurated 256$\times$256 DiT-L/2 samples.
+Figure J: Uncurated 256$\times$256 DiT-L/2 samples.
 Classifier-free guidance scale = 2.0.
 Class label = “lake shore" (975)
-Figure I: Uncurated 256$\times$256 DiT-L/2 samples.
+Figure K: Uncurated 256$\times$256 DiT-L/2 samples.
 Classifier-free guidance scale = 2.0.
 Class label = “lion" (291)
-Appendix H Broader Impacts
+Appendix H Discussion on Similarity with the Previous Work
+While both our work and (Go et al., 2023a) explore the characteristics of denoising tasks in diffusion models, the aspects of exploration in each work are substantially different.
+The notion of task affinity introduced in (Fifty et al., 2021, Go et al., 2023a) refers to how harmoniously the model can learn multiple tasks together.
+Specifically, their work focuses on identifying and mitigating conflicts between tasks, emphasizing task interactions and transferability by analyzing task similarities (e.g., gradient similarity or alignment).
+In contrast, our work explicitly quantifies the relative difficulty of individual denoising tasks across timesteps as a standalone property, independent of task interdependencies.
+The analysis of task difficulty in our work involves evaluating metrics such as loss behavior or convergence rates, directly reflecting the complexity of solving each task at different timesteps.
+Therefore, while (Go et al., 2023a) addresses how tasks relate and interact during multi-task learning, our focus lies in systematically characterizing the intrinsic difficulty of tasks across timesteps in diffusion models.
+Appendix I Broader Impacts
 Generative models, such as diffusion models, have the potential to significantly impact society, particularly through DeepFake applications and the use of biased datasets.
@@
 This situation may result in the dissemination of incorrect information and the manipulation of public opinion.
-Appendix I Limitations
+Appendix J Limitations
 In this work, we demonstrated the varying difficulties of denoising tasks through empirical results on various diffusion frameworks and proposed a curriculum learning approach that effectively enhances diffusion model training.